### v1

# Resonant Photon-Axion Mixing Driven by Dark Matter Oscillations

Abstract
Wave propagation in a coherently oscillating background is intrinsically a periodically driven problem. We show that photon propagation through an axion dark matter background in the presence of a magnetic field is governed by Floquet physics, distinct from conventional static or adiabatic mixing paradigms. Coherent photon–axion mode conversion occurs when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$) with the axion mass $m_{a}$, even far from the standard level-crossing condition $\Delta_{\gamma}\approx\Delta_{a}$. Crucially, this resonance disappears entirely if the axion oscillations are averaged over, and is therefore systematically missed in conventional static or adiabatic treatments. This driven resonance represents a unitary Floquet mode-mixing process and is fundamentally distinct from parametric instability or stimulated axion decay, preserving the axion dark matter number density. We develop a general Floquet framework for photon propagation in oscillating axion backgrounds, revealing that resonant mixing generates robust polarization effects during propagation. As an astrophysical application, we apply this mechanism to the realistic environment of the blazar 3C 279 to derive concrete constraints on the axion–photon coupling. While the observational manifestation depends on environmental conditions, the underlying driven mixing mechanism is generic to coherent axion dark matter, revealing a previously overlooked regime of photon–axion conversion.
Introduction.– The propagation of waves in time-dependent media is a central problem across physics. When the properties of a medium vary periodically in time, the resulting dynamics are generically governed by Floquet physics, leading to resonant mode mixing that cannot be captured by static or adiabatic approximations. Such driven propagation phenomena are well known in condensed matter and quantum optics [1965PhRv..138..979S, Sambe:1973cnm, 2004PhR...390....1C, Oka:2009djg, Eckardt:2016lof], yet they have received little attention in the context of particle astrophysics.
A prominent and well-motivated realization of a coherently oscillating medium arises if dark matter (DM) [Zwicky:1933gu, Bertone:2018krk] is composed of axions [Peccei:1977hh, Peccei:1977ur, Weinberg:1977ma, Wilczek:1977pj, Cheng:1987gp] or axion-like particles [Preskill:1982cy, Abbott:1982af, Dine:1982ah, Arvanitaki:2009fg, Graham:2015ouw, Peccei:2006as, Marsh:2015xka, DiLuzio:2020wdo, Adams:2022pbo, OHare:2024nmr]. In this case, the DM forms a classical field that oscillates at a frequency set by the axion mass over macroscopic coherence scales. Photons propagating through this background therefore experience a medium whose optical properties are intrinsically periodic in time. Nevertheless, conventional treatments of photon–axion mixing [Raffelt:1987im, Raffelt:1996wa] typically average over the axion oscillations or assume an effectively static background, thereby reducing the problem to a time-independent level-crossing or adiabatic conversion scenario.
In astrophysical environments, the photon–axion system is typically analyzed in two regimes: photon–axion conversion (the Primakoff effect) [Primakoff:1951iae, Sikivie:1983ip, Raffelt:1987im] and rotation of the polarization plane (birefringence) [Carroll:1989vb, Carroll:1998zi, Harari:1992ea]. Conventional analyses employ adiabatic or perturbative approximations, effectively averaging over rapid axion oscillations or treating the background as quasi-static.
In this work, we show that these assumptions miss a qualitatively novel class of resonant propagation effects. We demonstrate that photon–axion mixing in a coherent axion DM background is fundamentally a periodically driven problem, governed by Floquet dynamics rather than static dispersion matching. Our analysis reveals that the intrinsic time periodicity of the axion background acts as a parametric driver, actively inducing coherent mixing between photon and axion modes. This process leads to parametric resonance when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion mass as $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$). Consequently, resonant conversion can occur far from the conventional level-crossing resonance condition $\Delta_{\gamma}\approx\Delta_{a}$ [Pshirkov:2007st, Hook:2018iia]. This resonance arises purely from temporal phase matching induced by the oscillating background and does not rely on spatial inhomogeneities or fine-tuned environmental profiles. The driven resonance is not a higher-order correction to static mixing: it relies on the coherent axion oscillation and vanishes completely once this time dependence is averaged out.
The driven resonance identified here is fundamentally distinct from parametric instability or stimulated axion decay ($m_{a}\simeq 2\omega$) [Tkachev:2014dpa, Arza:2019nta, Caputo:2018vmy, Caputo:2018ljp]. The propagation dynamics remain unitary, the axion background number density is conserved, and the system exhibits coherent Floquet mode mixing rather than exponential amplification.
We develop a general Floquet framework for photon propagation in oscillating axion backgrounds and show that this mechanism leads to robust and generic consequences for photon polarization during propagation. While specific observational manifestations depend on environmental conditions, the underlying driven mixing phenomenon is intrinsic to coherent axion DM and represents a previously overlooked regime of photon–axion conversion. Our results establish a systematic framework for analyzing propagation effects in time-dependent backgrounds, revealing previously unexplored resonant regimes.
Driven photon–axion Mixing.– Consider photon propagation in the presence of a magnetic field and a coherently oscillating axion background (see also [Espriu:2014lma, Masaki:2019ggg, McDonald:2019wou] for a similar system),
$\bar{a}(t,z)=m_{a}^{-1}\sqrt{2\rho_{\text{DM}}}\cos(m_{a}t+\varphi_{a})$ (1)
where $m_{a}$ is the axion mass, $\rho_{\text{DM}}$ is the DM density, and $\varphi_{a}$ is the phase of the axion background. In an appropriate polarization basis, the propagation equation takes the form
$i\partial_{z}\Psi=\left[\mathcal{H}_{0}+\mathcal{H}_{1}\right]\Psi$ (2)
where the state vector $\Psi$ includes the two photon polarization states and the axion field. The photon–axion system is governed by a Hermitian mixing Hamiltonian whose coefficients contain an explicit periodic modulation at frequency $m_{a}$, induced by the axion background. $\mathcal{H}_{0}$ incorporates the usual photon dispersion $\Delta_{\perp}$ and $\Delta_{\parallel}$, axion mass term $\Delta_{a}$, and magnetic-field-induced mixing $\Delta_{\mathrm{B}}$, and is given by
$\mathcal{H}_{0}=-\begin{pmatrix}\Delta_{\perp}&0&0\\ 0&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$. (3)
$\mathcal{H}_{1}$ encodes the axion-induced birefringence
$\mathcal{H}_{1}(z)=-\begin{pmatrix}0&\Delta_{F}&0\\ \Delta_{F}^{*}&0&0\\ 0&0&0\end{pmatrix}$, (4)
with
$\Delta_{F}=ig_{a\gamma}(\partial_{t}{\bar{a}}+\partial_{z}\bar{a})/2.$ (5)
In the linear polarization basis, the axion-induced Faraday term is purely imaginary, reflecting its interpretation as a rotation generator. The Hamiltonian is manifestly Hermitian, and its only explicit time dependence arises from the oscillating axion background:
$\Delta_{F}=-i\Delta_{\text{DM}}\sin(m_{a}t+\varphi_{a}),$ (6)
where the amplitude is $\Delta_{\text{DM}}\equiv g_{a\gamma}\sqrt{\rho_{\text{DM}}/2}$. Along the photon trajectory we take the eikonal propagation approximation $t\simeq z$ (with $c=1$).
The propagation equation represents a three-level driven system with period $T=2\pi/m_{a}$. It is convenient to remove the explicit time dependence induced by the axion background by transforming to a co-rotating frame that follows the axion-induced phase modulation. This is achieved by a time-dependent unitary transformation that eliminates the oscillatory off-diagonal term at the expense of introducing harmonic couplings between propagation eigenstates.
We perform a time-dependent unitary transformation $\Psi(z)=U(z)\Phi(z)$ to reallocate the rapid oscillations. Applying the rotation operator $U=u\oplus 1$, where $u$ acts on the photon subspace as
$u(z)=\exp\left[-i\theta(z)\sigma_{y}\right],$ (7)
the fast oscillating sinusoidal term is absorbed into a phase-dependent rotation angle $\theta(z)=\alpha\cos\phi(z)$ with $\alpha=\Delta_{\text{DM}}/m_{a}$ and $\phi(z)=m_{a}z+\varphi_{a}$. This rotation is chosen such that the term generated by $U^{\dagger}i\partial_{z}U$ exactly cancels the rapidly oscillating birefringence $\Delta_{F}$. Assuming negligible QED birefringence with $|\Delta_{\perp}-\Delta_{\parallel}|/2\ll m_{a}$ and $\Delta_{\mathrm{DM}}$, the effective Hamiltonian becomes:
$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}&0&\Delta_{\mathrm{B}}\sin\theta\\ 0&\Delta_{\gamma}&\Delta_{\mathrm{B}}\cos\theta\\ \Delta_{\mathrm{B}}\sin\theta&\Delta_{\mathrm{B}}\cos\theta&\Delta_{a}\end{pmatrix}$, (8)
where $\Delta_{\gamma}=(\Delta_{\parallel}+\Delta_{\perp})/2$ is the mean photon refractive index. The unitary rotation is exact algebraically and does not change the quasienergy spectrum; it simply relocates the periodic dependence.
In the regime $\Delta_{\mathrm{B}}\ll|\Delta_{\gamma}-\Delta_{a}|$, the off-resonant photon mode can be adiabatically eliminated, allowing the dynamics to be analyzed by treating each photon mode-axion pair independently within the rotating-wave approximation. The resulting two-level systems are governed by the couplings $G^{(1)}=\Delta_{\mathrm{B}}\sin\theta$ and $G^{(2)}=\Delta_{\mathrm{B}}\cos\theta$. These couplings depend on $\theta(z)$; $\sin\theta$ and $\cos\theta$ can be written as linear combinations of $e^{\pm i\theta}$.
Using the Jacobi-Anger expansion, $e^{i\alpha\cos\phi}=\sum_{\ell}i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi}$ with the Bessel function $J_{\ell}(\alpha)$, we expand $G^{(1,2)}$ into Fourier harmonics. The couplings can be written as
$G(z)=\sum_{\ell\in\mathbb{Z}}g_{\ell}e^{i\ell m_{a}z},$ (9)
where $g_{\ell}^{(1)}$ is nonzero only for odd $\ell$, while $g_{\ell}^{(2)}$ is nonzero only for even $\ell$. Explicitly, the $\ell$-th harmonic coupling scales as $g_{\ell}\sim\Delta_{\mathrm{B}}J_{\ell}(\alpha)$ up to phase factors, where the Bessel function $J_{\ell}(\alpha)$ behaves roughly as an oscillatory sine or cosine that decays proportionally to $\alpha^{-1/2}$.
In this co-rotating basis, the driven photon–axion system acquires a Floquet structure: the axion oscillation generates sidebands separated by integer multiples of $m_{a}$, and resonant mode mixing occurs when the propagation eigenvalue splitting matches one of these harmonics. In this basis, the axion-induced term acts as a periodic off-diagonal drive that couples propagation eigenstates separated in quasi-energy. Transitions between eigenstates become resonant whenever the energy mismatch $\Delta=\Delta_{\gamma}-\Delta_{a}$ is compensated by the absorption or emission of integer multiples of the driving frequency.
We remove the free propagation phases by defining $\Phi=\operatorname{diag}\left(e^{i\Delta_{\gamma}z},e^{i\Delta_{a}z}\right)\tilde{\Phi}_{\gamma}$. Under this transformation, the $\ell$-th Fourier component of the coupling acquires a phase factor $e^{i(\Delta_{\gamma}-\Delta_{a}-\ell m_{a})z}$, with the corresponding detuning given by $\Delta_{\ell}=\Delta_{\gamma}-\Delta_{a}-\ell m_{a}$. Resonance occurs when the phase is stationary, yielding the Floquet resonance condition:
$\Delta_{\gamma}-\Delta_{a}\approx nm_{a},\quad n\in\mathbb{Z}.$ (10)
This condition characterizes coherent mode mixing within a time-periodic background, demonstrating that substantial conversion can occur even far from the standard level-crossing resonance $\Delta_{\gamma}=\Delta_{a}$.
Near a given resonance $n$, keeping only the corresponding Fourier mode amounts to a rotating-wave approximation. The two-state system undergoes Rabi oscillations with an effective frequency $\Omega_{\rm R}=\sqrt{|g_{n}|^{2}+(\Delta_{n}/2)^{2}}$. The conversion probability is given by
$P_{\gamma\to a}\approx\frac{|g_{n}|^{2}}{\Omega_{\rm R}^{2}}\sin^{2}\big(\Omega_{\rm R}\,z\big).$ (11)
The resonance is maintained as long as the detuning satisfies $|\Delta_{n}(z)|\lesssim 2|g_{n}|$, a regime where the peak conversion efficiency remains at least 50%. Note that the photon state $\gamma$ here refers to the effective photon mode under rotation. It can be verified that in the static limit $m_{a}\to 0$, all sidebands collapse and the system reduces exactly to conventional photon–axion mixing. Additionally, for small modulation amplitude $\Delta_{\mathrm{DM}}\ll m_{a}$, only the lowest harmonics contribute, and the effect reduces to a single-frequency perturbation of the standard oscillation probability.
More detailed derivations are provided in the supplementary material. An equivalent description in the circular polarization basis is also discussed in the supplementary material, where the axion background appears as a periodic modulation of the refractive indices. Related effects of axion-induced frequency modulation have been discussed previously, primarily in the context of birefringence or modifications of the dispersion relation [McDonald:2019wou]. In those treatments, the time-dependent phase is typically treated perturbatively or averaged over, and the associated Floquet sidebands are not explicitly identified. As a result, the resonant mode mixing and the corresponding Rabi oscillations discussed here were not recognized.
We emphasize that the Floquet resonance described here is fundamentally different from parametric instability or stimulated axion decay with $m_{a}\simeq 2\omega$. Here, the DM number density remains conserved, with the oscillating background providing the necessary energy ”kick” to bridge the momentum mismatch between the photon and axion dispersion relations. The Floquet coherent mixing in a periodic background is also dynamically distinct from the phenomenon of parametric instability. While both involve frequency matching, the former constitutes a unitary driven two-level problem characterized by stable mode-switching. In contrast, parametric instability requires a non-Hermitian structure for energy pumping (typically associated with sum-frequency channels  [^1] In the context of stimulated axion decay, the instability of the photon–axion system—considering both axion DM and magnetic fields but omitting the plasma background—is explored in Ref. [Masaki:2019ggg], where an instability condition $\sqrt{k_{\gamma}^{2}+m_{a}^{2}}+k_{\gamma}=nm_{a}$ is empirically identified.), which is kinematically inaccessible in the difference-frequency channel. Consequently, our resonance analysis leads to coherent oscillations that preserve the background number density, rather than the exponential growth characteristic of unstable systems. More discussion of this distinction is provided in the supplementary material.
Representative Propagation.– To illustrate the physical implications of the Floquet resonance, we consider a representative photon propagation scenario in which the photon dispersion $\Delta_{\gamma}$ varies slowly along the propagation direction, while the axion background remains coherent over the relevant length scale. This setup captures the essential physics of resonance crossing without requiring detailed modeling of a specific astrophysical environment.
As the photons propagate, the slowly varying dispersion causes the mismatch $\Delta(z)=\Delta_{\gamma}(z)-\Delta_{a}$ to sweep across the Floquet resonance condition $\Delta=nm_{a}$. When this occurs, the driven mode coupling becomes resonant, leading to efficient photon–axion conversion. Importantly, this resonance depends on temporal phase matching rather than spatial level crossing, and therefore persists in the absence of sharp gradients.
Figure 1: Evolution of the photon–axion conversion probability $P_{\gamma\to a}$ along the jet of blazar. The blue solid line represents the case with an axion DM background, while the orange line shows the standard adiabatic conversion without DM. The vertical red dashed lines indicate the locations where the resonance condition $\Delta_{\gamma}(r)-\Delta_{a}\approx nm_{a}$ is satisfied.
Fig. 1 shows the photon–axion conversion probability $P_{\gamma\to a}$ as a function of propagation distance for a representative choice of parameters satisfying the Floquet resonance condition. A pronounced enhancement occurs when $\Delta(z)$ crosses $\Delta=m_{a}$, demonstrating coherent mode mixing induced by the axion-driven modulation. Away from resonance, the conversion probability remains small and oscillatory, consistent with conventional photon–axion mixing.
The resonant region acts as a phase-matched mixing layer, in which the axion oscillation continuously supplies the energy mismatch required for conversion. The resulting behavior is analogous to Rabi oscillations in driven two-level systems, with the conversion probability bounded by unitarity.
We have verified that the resonance persists under moderate variations of the dispersion profile and in the presence of stochastic fluctuations, provided that the axion field remains coherent over the resonance region. Thus, the effect is not a fine-tuned artifact of a specific background profile. Detailed numerical analysis is presented in the Supplementary Material.
A distinctive consequence of this driven mixing is the generation of polarization asymmetries between propagation modes. In particular, the stochastic crossing of Floquet resonances induces circular polarization with a characteristic variance set by the resonance strength. Since the axion-induced modulation affects the two circular polarization states with opposite signs, Floquet resonances generically break the symmetry between them. As a result, photon propagation through an axion DM background becomes intrinsically polarization-sensitive, even for initially unpolarized states.
In particular, repeated or stochastic crossings of Floquet resonances lead to a net circular polarization with vanishing mean but nonzero variance. This stochastic circular polarization arises from coherent, unitary mode mixing and is therefore qualitatively distinct from polarization generated by absorption or scattering processes. The magnitude of the variance is directly controlled by the resonance strength, thereby providing a direct probe of the driven mixing mechanism. Importantly, this polarization signature does not rely on fine-tuned source properties or detailed environmental modeling. It follows directly from the existence of Floquet resonances in photon propagation through a coherent axion background.
Constraints from the balazar 3C 279.– We validate this theoretical framework with full numerical solutions and apply it to the optical polarimetry of the blazar 3C 279 as an example. The parametric resonance breaks the symmetry between circular polarization states, inducing a stochastic signal in the blazar jet. Detailed configurations are provided in the supplementary material. To derive constraints on the axion–photon coupling $g_{a\gamma}$, we adopt a conservative statistical criterion: the standard deviation of the axion-induced circular polarization $\sigma_{\mathrm{std}}$ must not exceed the characteristic observational uncertainty of the blazar 3C 279 ($\sigma_{\Pi_{C}}\approx 0.30\%$ [Liodakis:2021els]). This ensures consistency between the predicted stochastic signal and the measured noise level. For a given astrophysical source, the axion phase $\varphi_{a}$ is effectively constant over observational timescales. However, since its value is a priori unknown, the axion-induced circular polarization behaves observationally as a stochastic contribution with zero mean. Even when the axion-induced circular polarization exhibits a bimodal distribution for certain masses, its sign and magnitude remain unpredictable due to the unknown axion phase, and the variance provides a conservative measure of its observational impact. Requiring its variance not to exceed the instrumental sensitivity therefore provides a conservative consistency bound. Fig. 2 shows the illustrative constraint on parameter space, with CAST bounds [CAST:2017uph] included for comparison.
For $m_{a}\in[10^{-23},10^{-20}]\,\mathrm{eV}$, we obtain conservative limits as $g_{a\gamma}\lesssim 2\times 10^{-11}\,\mathrm{GeV}^{-1}$. The distinctive band-like structure of the exclusion contour is a signature of the Floquet parametric resonance mechanism, reflecting the discrete nature of the resonance condition. Two features emerge: (1) The conversion efficiency, affected by the Bessel functions $J_{n}(\Delta_{\text{DM}}/m_{a})$, is maximized when the effective momentum mismatch is bridged by integer harmonics of $m_{a}$. The sharp dips in the exclusion limit correspond to ”sweet spots” where this resonance condition is optimally satisfied within the jet profile. (2) As $m_{a}$ varies, the dominant contribution shifts between different resonance orders $n$. Weaker constraints between bands reflect off-resonance regimes, where the system approaches the adiabatic limit, resulting in reduced circular polarization production.
The constraints in Fig. 2 show that axion-induced parametric resonance leads to verifiable physical consequences rather than negligible perturbative effects. Although the current limit region is set by the specific noise floor of 3C 279, the concrete limit validates the physical relevance of our framework.
Figure 2: Exclusion contours on the axion–photon coupling $g_{a\gamma}$ derived from 3C 279 optical polarimetry. The shaded region indicates parameters excluded by the non-observation of excess circular polarization. The resulting exclusion exhibits a distinctive band structure that directly reflects the discrete Floquet resonance condition $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, providing a characteristic signature of the driven mixing mechanism.
Conclusions.– Wave propagation in a coherently oscillating background is intrinsically a periodically driven problem. We have shown that photon propagation in axion DM therefore exhibits Floquet dynamics, leading to resonant photon–axion mode conversion beyond the conventional static or adiabatic mixing paradigm. This driven resonance arises when the dispersion mismatch between photons and axions is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, enabling coherent conversion even far from the standard level-crossing condition.
The resulting Floquet resonance represents a distinct class of propagation phenomena. Unlike traditional resonant conversion, which is localized to specific spatial regions where plasma and axion masses coincide, the driven resonance identified here originates from temporal phase matching induced by the oscillating DM background. As a consequence, resonant mixing can persist over extended propagation distances, allowing small conversion amplitudes to accumulate coherently and imprint robust signatures on photon polarization.
The significance of this mechanism depends on the stability of the driven system against environmental fluctuations. In rapidly varying media, strong gradients can disrupt the phase coherence required for Floquet mixing, suppressing the resonance. In sufficiently smooth environments, however, the driven propagation effect can dominate photon–axion conversion, opening a coherence-based window for probing axion dark matter that is complementary to localized, high-density resonance searches.
More broadly, the Floquet framework developed here is not restricted to a single axion species. Because the resonance condition relies on the periodic structure of the background rather than its microscopic identity, analogous driven conversion phenomena may arise in multi-axion systems or more complex dark-sector environments. Our results therefore establish a general foundation for studying resonant wave propagation in coherently oscillating media, revealing a previously unexplored regime of dark-matter-induced photon conversion.
Acknowledgements.
Acknowledgments.– The authors would like to thank the computing cluster of School of Fundamental Physics and Mathematical Sciences at Hangzhou Institute for Advanced Study, UCAS for providing the computational resources used in this work. This work is supported by the National Key Research and Development Program of China Grant No. 2020YFC2201502, and the National Natural Science Foundation of China under Grants No. 12475065, 12250010, 12447131, 12447105, and 12575113.
Supplementary Material for Resonant photon–axion Mixing
Driven by Dark Matter Oscillations
Run-Min Yao, Xiao-Jun Bi, Peng-Fei Yin, and Qing-Guo Huang
This Supplementary Material (SM) contains additional calculation and derivation in support of the results presented in this work. First, we give a full derivation of the mixing matrix $\mathcal{M}$. Then, we present additional details regarding the derivation of the two-level systems. Next, we discuss the relation between our work and parametric instability. Finally, we present a numerical analysis and apply the formalism to the relativistic jet of blazar 3C 279.
I photon–axion Conversion in an Axion Background
The Lagrangian for the photon field $A_{\mu}$ and axion field $a$ can be written as
$\mathcal{L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-A_{\mu}j^{\mu}+\frac{1}{2}(\partial_{\mu}a)^{2}-\frac{1}{2}m_{a}^{2}a^{2}-\frac{1}{4}g_{a\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu},$ (S1)
where $F_{\mu\nu}$ and $\tilde{F}^{\mu\nu}$ denote the electromagnetic field tensor and its dual tensor, respectively, $j^{\mu}=(\rho,\mathbf{J})$ is the electromagnetic current density, and $g_{a\gamma}$ is the axion–photon coupling. Adopting the temporal gauge $A_{0}=A^{0}=0$ and fixing the residual gauge freedom by imposing the Coulomb (radiation) gauge $\bm{\nabla}\cdot\mathbf{A}=0$, the equations of motion are:
$\partial_{t}^{2}\mathbf{A}-\nabla^{2}\mathbf{A}=\mathbf{J}+g_{a\gamma}\partial_{t}a\bm{\nabla}\times\mathbf{A}-g_{a\gamma}\bm{\nabla}a\times\partial_{t}\mathbf{A},$ (S2)
$-\bm{\nabla}\cdot\partial_{t}\mathbf{A}=\rho-g_{a\gamma}\bm{\nabla}a\cdot\left(\bm{\nabla}\times\mathbf{A}\right),$ (S3)
$\partial_{t}^{2}a-\nabla^{2}a+m_{a}^{2}a=-g_{a\gamma}\partial_{t}\mathbf{A}\cdot\left(\bm{\nabla}\times\mathbf{A}\right).$ (S4)
For a weakly magnetized cold plasma, the dielectric tensor is isotropic, $\bm{\epsilon}=(1-\frac{\omega_{\mathrm{pl}}^{2}}{\omega^{2}})\mathbb{I}$, where $\omega_{\mathrm{pl}}=\sqrt{4\pi\alpha n_{e}/m_{e}}$ is the plasma frequency and $\omega$ is the energy of the photon–axion system. In the radiation gauge, longitudinal plasma oscillations decouple from the propagating electromagnetic modes. In Fourier space, the induced plasma current therefore contributes an isotropic dispersion term $\mathbf{J}_{\mathrm{ind}}(\omega)=-\omega^{2}(\mathbb{I}-\bm{\epsilon})\mathbf{A}(\omega)=-\omega_{\mathrm{pl}}^{2}\mathbf{A}(\omega)$ to both photon polarizations, which can be absorbed into an effective photon mass.
We apply the linearized perturbation to the equation and write the fields into background and perturbation
$\mathbf{A}=\bar{\mathbf{A}}+\delta\mathbf{A},\qquad a=\bar{a}+\delta a,$ (S5)
where $\bar{a}$ and $\bar{\mathbf{B}}=\bm{\nabla}\times\bar{\mathbf{A}}$ are the background axion and magnetic fields. The perturbations satisfy the equations of motion:
$\partial_{t}^{2}\delta\mathbf{A}-\nabla^{2}\delta\mathbf{A}+\omega_{\mathrm{pl}}^{2}\delta\mathbf{A}=g_{a\gamma}\bar{\mathbf{B}}\partial_{t}\delta a+g_{a\gamma}\partial_{t}\bar{a}\bm{\nabla}\times\delta\mathbf{A}-g_{a\gamma}\bm{\nabla}\bar{a}\times\partial_{t}\delta\mathbf{A},$ (S6)
$-\bm{\nabla}\cdot\partial_{t}\delta\mathbf{A}=\delta\rho-g_{a\gamma}\bm{\nabla}\bar{a}\cdot\left(\bm{\nabla}\times\delta\mathbf{A}\right)-g_{a\gamma}\bm{\nabla}\delta a\cdot\bar{\mathbf{B}},$ (S7)
$\partial_{t}^{2}\delta a-\nabla^{2}\delta a+m_{a}^{2}\delta a=-g_{a\gamma}\bar{\mathbf{B}}\cdot\partial_{t}\delta\mathbf{A}.$ (S8)
For a magnetic field perpendicular to the photon propagation direction and an axion gradient parallel to it, the source terms in Eq.(S7) vanish. The last two terms in Eq. (S6) characterize the birefringence effect induced by the axion DM. The first terms on the RHS of Eq. (S6) and Eq. (S8) reflect the axion–photon couplings. In conventional analysis regarding the photon–axion conversion, the variation of the axion DM is neglected.
We choose the direction of the wave vector $\hat{\mathbf{k}}$ along the $z$-axis and adopt the plane-wave Ansatz
$\delta\mathbf{A}=-i\tilde{\mathbf{A}}(z)e^{-i\omega t},\qquad\delta a=\tilde{a}(z)e^{-i\omega t},$ (S9)
where $\delta\mathbf{A}$ includes an additional factor of i-i compared to $\delta a$. This choice ensure that the axion–photon coupling terms in the mixing matrix remain real, consistent with the standard form of the equations of motion found in most literature. Subsequently, we derive
[…]
$(\partial_{z}^{2}+\omega^{2}-m_{a}^{2})\tilde{a}=-g_{a\gamma}\bar{\mathbf{B}}\omega\tilde{\mathbf{A}}$. (S11)
The equations can be rearranged as
$(\partial_{z}^{2}+\omega^{2}+2\omega\mathcal{M})\psi=0,$ (S12)
[…]
$\mathcal{M}(z)=\begin{pmatrix}\Delta_{\perp}&\Delta_{F}&0\\ \Delta_{F}^{*}&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$, (S13)
where $\Delta_{\perp}=\Delta_{\parallel}\equiv-\omega_{\mathrm{pl}}^{2}/(2\omega)$, $\Delta_{a}\equiv-m_{a}^{2}/(2\omega)$, and $\Delta_{\mathrm{B}}\equiv g_{a\gamma}B/2$. The $\Delta_{F}$ term originates from the axion-gradient-induced birefringence and is formally analogous to the Faraday rotation term. Explicitly, it is given by
$\Delta_{F}\equiv\frac{g_{a\gamma}}{2\omega}(\partial_{t}\bar{a}\partial_{z}+i\omega\partial_{z}\bar{a})$ (S14)
where the operator $\partial_{z}$ in the first term acts on the photon field. Within the WKB approximation, $i\partial_{z}\psi\approx\omega\psi$, the operator $\partial_{z}$ acting on $\psi$ can be replaced by $i\omega$, yielding
$\Delta_{F}=\frac{i}{2}g_{a\gamma}(\partial_{t}\bar{a}+\partial_{z}\bar{a}).$ (S15)
The second-order equations of motion can be expressed in first-order form
$(i\partial_{z}+\omega+\mathcal{M})\psi=0.$ (S16)
In the mixing matrix, we have neglected the absorption effects of photons, the contribution of the cosmic microwave background’s energy density to the dispersion relation, and the QED vacuum birefringence effect.
Assuming a small DM virial velocity $v_{\mathrm{DM}}\ll 1$, we model the background field as
$\bar{a}(t,z)=m_{a}^{-1}\sqrt{2\rho_{\text{DM}}}\cos(m_{a}t+\varphi_{a}),$ (S17)
where $\rho_{\text{DM}}$ is the DM density and $\varphi_{a}$ is the phase of the axion background. This leads to an oscillating birefringence term:
$\Delta_{F}=-i\Delta_{\text{DM}}\sin(m_{a}t+\varphi_{a}^{\prime}),$ (S18)
where the amplitude is $\Delta_{\text{DM}}\equiv g_{a\gamma}\sqrt{(1+\mathcal{F}^{2})\rho_{\text{DM}}/2}$ and the phase is $\varphi_{a}^{\prime}=\varphi_{a}+\arctan\mathcal{F}$. The factor $\mathcal{F}(z)=m_{a}^{-1}d\ln\rho_{\text{DM}}/dz$ can be used to describe possible corrections from spatial DM gradients, but will be neglected in the following. For fuzzy DM, $\mathcal{F}(z)$ is less than the virial velocity of DM. Along the photon trajectory we take the eikonal propagation approximation $t\simeq z$ (with $c=1$).
II Effective Two-Level systems
Applying $\psi(z)=e^{i\omega z}\Psi(z)$, the reduced Schrödinger-like equation can be written as
$i\frac{d}{dz}\Psi(z)=\mathcal{H}(z)\Psi(z),\quad\mathcal{H}(z)=-\mathcal{M}(z),$ (S19)
where $\mathcal{M}$ is the mixing matrix including the oscillating birefringence term. To solve the dynamics analytically, we perform a time-dependent unitary transformation $\Psi(z)=U(z)\Phi(z)$ to eliminate the rapid oscillations in the photon sub-block. The effective Hamiltonian becomes
$\mathcal{H}^{\prime}(z)=U^{\dagger}\mathcal{H}U-iU^{\dagger}\frac{dU}{dz}.$ (S20)
Define $\phi(z)\equiv m_{a}z+\varphi_{a}$, the photon sub-sector of $\mathcal{H}$ can be decomposed using Pauli matrices:
$\mathcal{H}_{\gamma}(z)=-\Delta_{\gamma}\,\mathbb{I}_{2}-\delta\Delta_{\gamma}\,\sigma_{z}-\Delta_{\text{DM}}\sin\phi\,\sigma_{y},$ (S21)
where $\Delta_{\gamma}=(\Delta_{\perp}+\Delta_{\parallel})/2$ and $\delta\Delta_{\gamma}=(\Delta_{\perp}-\Delta_{\parallel})/2$. We introduce a time-dependent unitary rotation acting in the photon subspace
$u(z)=\exp\!\big[-i\theta(z)\sigma_{y}\big],\quad\theta(z)=\alpha\cos\phi(z).$ (S22)
The transformation matrix is
$U(z)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}$. (S23)
The term $iu^{\dagger}\partial_{z}u$ can be calculated as
$iu^{\dagger}\partial_{z}u=-\alpha m_{a}\sin\phi\;u^{\dagger}\sigma_{y}u=-\alpha m_{a}\sin\phi\;\sigma_{y},$ (S24)
where the final step utilises the fact that $\sigma_{y}$ commutes with itself. To cancel the $\Delta_{\text{DM}}\sin\phi\,\sigma_{y}$ term in the Hamiltonian, we set $\alpha=\Delta_{\text{DM}}/m_{a}$, yielding the effective Hamiltonian
$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}+\delta\Delta_{\gamma}\cos(2\theta)&\delta\Delta_{\gamma}\sin(2\theta)&\Delta_{B}\sin\theta\\[6.0pt] \delta\Delta_{\gamma}\sin(2\theta)&\Delta_{\gamma}-\delta\Delta_{\gamma}\cos(2\theta)&\Delta_{B}\cos\theta\\[6.0pt] \Delta_{B}\sin\theta&\Delta_{B}\cos\theta&\Delta_{a}\end{pmatrix}$. (S25)
When $|\delta\Delta_{\gamma}|\ll m_{a}$ and $\Delta_{\mathrm{DM}}$, the phase evolution rate due to QED birefringence becomes negligible compared to the DM modulation frequency. In this regime, the photon states are effectively degenerate for axions, allowing the QED contribution to be safely neglected. The transformed Hamiltonian simplifies to
$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}&0&\Delta_{\mathrm{B}}\sin\theta\\ 0&\Delta_{\gamma}&\Delta_{\mathrm{B}}\cos\theta\\ \Delta_{\mathrm{B}}\sin\theta&\Delta_{\mathrm{B}}\cos\theta&\Delta_{a}\end{pmatrix}$. (S26)
For $\Delta_{\mathrm{B}}\ll|\Delta_{\gamma}-\Delta_{a}|$, the dynamics can be analyzed by treating each photon mode-axion pair independently under the rotating-wave approximation. The two-state systems are governed by couplings $G^{(1)}=\Delta_{\mathrm{B}}\sin\theta$ and $G^{(2)}=\Delta_{\mathrm{B}}\cos\theta$. Expressing $\sin\theta$ and $\cos\theta$ via the Jacobi–Anger identity
$e^{i\alpha\cos\phi}=\sum_{\ell=-\infty}^{\infty}i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi},$ (S27)
we expand $G^{(1,2)}$ into Fourier harmonics:
$G^{(1)}(z)$ $=\frac{\Delta_{\mathrm{B}}}{2i}\sum_{\ell}[1-(-1)^{\ell}]i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi},$ (S28)
$G^{(2)}(z)$ $=\frac{\Delta_{\mathrm{B}}}{2}\sum_{\ell}[1+(-1)^{\ell}]i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi}.$ (S29)
Thus, each photon mode couples to the axion through a series of sidebands with frequencies $\ell m_{a}$. $G^{(1)}$ and $G^{(2)}$ carry odd and even harmonics, respectively. We define the relevant coupling as
$G(z)=\sum_{\ell\in\mathbb{Z}}g_{\ell}e^{i\ell m_{a}z},$ (S30)
with
$g_{\ell}^{(1)}$ $=\frac{[1-(-1)^{\ell}]i^{\ell-1}e^{i\ell\varphi_{a}}}{2}J_{\ell}(\alpha)\Delta_{\mathrm{B}},$ (S31)
$g_{\ell}^{(2)}$ $=\frac{[1+(-1)^{\ell}]i^{\ell}e^{i\ell\varphi_{a}}}{2}J_{\ell}(\alpha)\Delta_{\mathrm{B}}.$ (S32)
Focusing on a single photon mode, the coupled equations become:
$-i\frac{d}{dz}\Phi_{\gamma}=\Delta_{\gamma}\Phi_{\gamma}+G(z)\Phi_{a},$ (S33)
$-i\frac{d}{dz}\Phi_{a}=G(z)\Phi_{\gamma}+\Delta_{a}\Phi_{a}.$ (S34)
Moving to the interaction picture via $\Phi_{\gamma}=\tilde{\Phi}_{\gamma}e^{i\Delta_{\gamma}z}$, $\Phi_{a}=\tilde{\Phi}_{a}e^{i\Delta_{a}z}$, we obtain
$-i\frac{d}{dz}\tilde{\Phi}_{\gamma}=\sum_{\ell}g_{\ell}e^{-i\Delta_{\ell}z}\tilde{\Phi}_{a},$ (S35)
$-i\frac{d}{dz}\tilde{\Phi}_{a}=\sum_{\ell}g_{\ell}^{*}e^{i\Delta_{\ell}z}\tilde{\Phi}_{\gamma},$ (S36)
with
$\Delta_{\ell}=\Delta_{\gamma}-\Delta_{a}-\ell m_{a}.$ (S37)
Resonance occurs when the phase oscillation is slow, i.e., when the momentum mismatch is compensated by the axion mass harmonics. Retaining the resonant harmonic $\ell=n$ that satisfies $\Delta_{n}\approx 0$, we obtain the Rabi oscillation dynamics with an effective Rabi frequency
$\Omega_{\rm R}=\sqrt{|g_{n}|^{2}+(\Delta_{n}/2)^{2}}.$ (S38)
The conversion probability is given by:
$P_{\gamma\to a}=\frac{|g_{n}|^{2}}{\Omega_{\rm R}^{2}}\sin^{2}\big(\Omega_{\rm R}\,z\big).$ (S39)
Additionally, we discuss the choice of transformations for diagonalizing the Hamiltonian $\mathcal{H}$. In general, Floquet theory can be directly applied to the three-level system. However, to isolate the dominant resonant dynamics and obtain a transparent physical picture, we diagonalize the photon subspace and reduce the problem to effective two-level photon–axion systems. An intuitive choice is to transform to the circular polarization basis using
$U_{C}(z)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i&0\\ 1&-i&0\\ 0&0&1\end{pmatrix}$. (S40)
The Hamiltonian is given by
$\mathcal{H}_{C}^{\prime}=-\begin{pmatrix}\Delta_{+}&\delta\Delta_{\gamma}&\frac{\Delta_{\mathrm{B}}}{\sqrt{2}}\\ \delta\Delta_{\gamma}&\Delta_{-}&\frac{i\Delta_{\mathrm{B}}}{\sqrt{2}}\\ \frac{\Delta_{\mathrm{B}}}{\sqrt{2}}&-\frac{i\Delta_{\mathrm{B}}}{\sqrt{2}}&\Delta_{a}\end{pmatrix}$ (S41)
where $\Delta_{\pm}=\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi$. Neglecting the QED birefringence effect, the evolution for the photon state can be formally written as
$\Phi_{\pm}$ $\simeq\exp\left[-i\int^{z}dz^{\prime}(\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi(z^{\prime}))\right]\tilde{\Phi}_{\pm}$ (S42)
$=e^{-i\Delta_{\gamma}z}e^{\pm i\alpha\cos\phi}\tilde{\Phi}_{\pm}$. (S43)
Repeating the Fourier harmonic expansion as above, one would also obtain the same resonance condition and Rabi oscillations. Although axion-induced frequency modulation has been noted in earlier studies [McDonald:2019wou], it is often incorporated as a slowly varying phase or an effective modification of the refractive index. Such treatments effectively average over the harmonic structure of the modulation and therefore do not capture the Floquet sidebands responsible for resonant mode mixing.
The transformations $U(z)$ and $U_{C}$ offer two distinct yet complementary physical interpretations. The transformation $U(z)$ represents a co-rotating frame perspective. By shifting into a basis that rotates alongside the axion-induced birefringence effect, we effectively eliminate the rapid oscillations of the off-diagonal mixing terms. In this frame, the interaction is understood through the lens of geometric phase, where the coupling to the axion arises from the mismatch between the photon’s polarization drift and the background field’s oscillation frequency.
The circular basis $U_{C}$ provides a frequency modulation perspective. Here, the axion field acts as a dynamic medium that oscillates the refractive indices of the left- and right-handed circular polarizations in opposite phases. The resonance condition is explained by the Jacobi-Anger expansion, where the carrier frequency of the photon acquires sidebands at integer multiples of the axion mass $m_{a}$. While both methods yield identical Rabi frequencies for the resonant transition, the circular basis is often more robust when incorporating higher-order corrections, such as QED birefringence, as it treats the two helicity states as the fundamental vacuum eigenmodes.
III Coherent Mode Mixing vs. Parametric Instability
In the literature on axion electrodynamics, it is well established that an oscillating axion background can induce a parametric instability in the electromagnetic field. This is often formulated using the Mathieu equation and interpreted as the stimulated decay of axions into photons. Superficially, this resembles the propagation equation, Eq. (S12), utilized in this work, where the mixing matrix $\mathcal{M}(z)$ also contains a periodic contribution from the axion DM background. In this appendix, we clarify that these two phenomena, despite originating from the same Lagrangian, describe physically distinct regimes and yield qualitatively different results.
III.1 Common Origin
Both formalisms derive from the standard axion–photon Lagrangian Eq. (S1). Consider a classical, coherently oscillating axion background
$a(t)\sim a_{0}\cos(m_{a}t).$ (S44)
Maxwell’s equations acquire an explicit time dependence, characterized by Eqs. (S6) to (S8). However, the subsequent reduction of these equations depends critically on the physical boundary conditions and the kinematic regime of interest.
III.2 Regime I: Parametric Instability (The Mathieu Limit)
The analysis of axion-induced parametric instability typically focuses on the spontaneous production and amplification of photon modes from the vacuum (or a low-occupation state). In this regime, one must retain the full second-order time dynamics, including both positive- and negative-frequency components of the field momentum modes:
$\mathbf{A}(\mathbf{x},t)\sim A_{\pm}(t)\,e^{i\mathbf{k}\cdot\mathbf{x}},\qquad A_{\pm}(t)=c(t)e^{-i\omega t}+c^{*}(t)e^{+i\omega t}.$ (S45)
For circular polarizations, the equation of motion reduces to a Mathieu equation:
$\ddot{A}_{\pm}+\omega^{2}\left[1\pm\epsilon\cos(m_{a}t)\right]A_{\pm}=0,$ (S46)
where $\epsilon\sim g_{a\gamma}a_{0}m_{a}/\omega$. The defining feature of this regime is that the periodic axion background modulates the effective frequency squared of the oscillator. This modulation couples the positive-frequency ($e^{-i\omega t}$) and negative-frequency ($e^{+i\omega t}$) modes. When $m_{a}\simeq 2\omega$, the system enters an instability band where the field amplitude grows exponentially. This corresponds to the stimulated decay of the axion field into photon pairs, a genuine parametric instability where the background acts as an energy pump.
III.3 Regime II: Coherent Mode Mixing (The Propagation Limit)
In contrast, this work investigates the propagation of a pre-existing electromagnetic wave with a fixed carrier frequency $\omega$ and well-defined direction. Employing the slowly varying envelope approximation, we express the field as a slowly evolving amplitude modulating a rapid oscillatory phase. Crucially, this procedure effectively integrates out backward-propagating and negative-frequency modes. The resulting dynamics are governed by Eq. (S12). While the axion background still introduces a periodic modulation in the mixing matrix $\mathcal{M}(z)$, its physical role changes fundamentally:
  • The modulation acts as a linear, Hermitian coupling between forward-propagating modes (e.g., photon polarization states).
  • It does not couple positive- and negative-frequency solutions.
Consequently, the system does not exhibit Mathieu-type exponential instabilities. Instead, it displays Floquet resonances—enhanced but bounded mode conversion—when the eigenvalue difference matches the modulation frequency: $\lambda_{i}-\lambda_{j}=nm_{a}$.
III.4 Theoretical Synthesis
The distinction between these two phenomena can be understood by analogy to quantum optical systems.
  • Parametric Instability (Regime I) is analogous to a driven harmonic oscillator. A periodic modulation of the oscillator’s parameters leads to exponential growth of the mode amplitude within instability bands. In quantum terms, this represents parametric amplification, or particle creation from the vacuum.
  • Floquet Mixing (Regime II) is analogous to a driven multi-level system, such as Rabi oscillations in atomic physics. The relevant degrees of freedom are discrete internal states, and the periodic background induces transitions among them. Probabilities oscillate but remain bounded (unitary evolution).
From a formal Floquet theory perspective, both systems involve linear differential equations with periodic coefficients. However, the spectral implications differ:
  • In the Mathieu case (Instability), the Floquet quasi-energies can acquire imaginary parts, signaling an unbounded instability.
  • In the Propagation case (Mixing), the quasi-energies remain real. Resonances here correspond to avoided crossings in the quasi-energy spectrum, maximizing the mixing angle between states rather than the total energy of the system.
Thus, while both phenomena arise within the framework of periodically driven systems, the propagation formalism employed in this work isolates the physics of coherent state mixing, explicitly distinct from the particle-production physics of parametric instability.
IV Numerical analysis
To numerically verify the analytical framework, we solve the Eq. (S16) incorporating the mixing matrix $\mathcal{M}$ for a concrete astrophysical environment. We model the relativistic jet of blazar 3C 279, focusing on the unpolarized optical R-band photons ($E_{\gamma}\approx 1.9\,\mathrm{eV}$) with a representative parameter set: an axion mass $m_{a}=10^{-22}\,\mathrm{eV}$ and a coupling constant $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$. The radiation zone, located $r_{0}=0.5\,\mathrm{pc}$ from the central engine, is characterized by a magnetic field $B_{0}=0.5\,\mathrm{G}$ and electron density $n_{e,0}=5\times 10^{3}\,\mathrm{cm}^{-3}$ [A:2025wjt]. The jet is modeled as a conical outflow with a Doppler factor $\delta\simeq 13$ [A:2025wjt], assuming scaling profiles $B(r)\propto r^{-1}$ and $n_{e}(r)\propto r^{-2}$ over a total length $L_{\rm jet}=1\,\mathrm{kpc}$. To capture the fine-grained evolution of the mixing process, the calculation is performed over a domain size ranging from $5\times 10^{-7}\,\mathrm{pc}$ to $10^{-3}\,\mathrm{pc}$. To isolate and illustrate the resonance structures dictated by the plasma profile, we adopt a uniform DM density $\rho_{\rm DM}=1\,\mathrm{GeV\cdot cm^{-3}}$ ($2.63\times 10^{7}\,M_{\odot}\mathrm{\cdot kpc^{-3}}$).
The numerical results of the photon–axion conversion probability in Fig. 1 demonstrate resonant conversion exactly at the locations predicted by Eq. (10), while the behavior reduces to the adiabatic limit away from resonance. This excellent agreement confirms that the Floquet analysis accurately describes the physics of the driven system. The conversion probability exhibits a distinct step-like growth, indicating that resonant conversion occurs within narrow spatial windows where the phase-matching condition is satisfied. The resonance positions depend solely on the jet’s plasma profile and the axion mass, independent of the DM density $\rho_{\rm DM}$. Thus, while a realistic non-uniform DM halo would modulate the amplitude of each ”step” according to local $\rho_{\rm DM}$ and $B$ values, the qualitative structure of multiple resonances remains unchanged.
The spatial width of the parametric resonance $L_{\text{res}}$ can be estimated as $\left|4g_{n}/(\nabla_{r}\Delta_{n})\right|_{r_{\text{res}}}$. For the jet model considered here, the resonance width is approximately $2(nm_{a})^{-1}g_{a\gamma}B_{0}J_{n}(\alpha)r_{0}$. For the parameters used in Fig. 1, we obtain $L_{\text{res}}\sim 3\,\text{pc}$, in agreement with the numerical results. This finite width confirms that the resonance is robust against density fluctuations on scales smaller than $L_{\text{res}}$. Given that $L_{\text{res}}$ is two orders of magnitude smaller than the total jet length ($\sim\mathrm{kpc}$), plasma density variations within each resonance region are negligible. This separation of scales ensures that the parametric resonance remains robust even in highly inhomogeneous jet environments, where the global density gradient merely serves to sequentially satisfy discrete resonance conditions.
The parametric resonance relies on the coherent oscillation of the axion field. The validity of this assumption depends on the coherence time of the DM field, $\tau_{\mathrm{coh}}\sim(m_{a}v^{2})^{-1}$, compared to the photon’s transit time through the resonance region, $t_{\mathrm{cross}}\sim L_{\mathrm{res}}$. For ultralight axions considered here with virial velocity $v\sim 10^{-3}$, the coherence time is $\tau_{\mathrm{coh}}\sim\mathcal{O}(10^{5})\,\text{years}$. In contrast, the spatial width $L_{\mathrm{res}}\sim 3\,\mathrm{pc}$ corresponds to a transit time $t_{\mathrm{cross}}\approx 10\,\text{years}$. The condition $t_{\mathrm{cross}}\ll\tau_{\mathrm{coh}}$ ensures that the DM background acts as a highly coherent driving force during the resonant interaction. Moreover, the fractional frequency spread $\delta m_{a}/m_{a}\sim 10^{-6}$ induces a negligible broadening of the resonance layer compared to its physical width determined by the jet geometry and coupling strength.
Figure S1: ECDF of the photon-to-axion conversion probability $P_{\gamma\to a}$. The color coding indicates various turbulence correlation lengths $L_{fluc}$ for electron density (from $10^{-3}$ to $1\text{ pc}$), while solid and dashed lines distinguish between models with and without DM background, respectively. Each curve is derived from a sample size of $N=1000$. Note that the x-axis is presented in log-scale.
To assess the robustness of the resonance mechanism, we introduce 50% stochastic electron density fluctuations across turbulence scales ranging from $L_{\rm fluc}=10^{-3}$ to $1$ pc. As shown in Fig. S1, the mean conversion probability in the DM background remains consistently 1-2 orders of magnitude higher than the standard baseline without DM, regardless of the correlation length. While small-scale fluctuations ($L_{\rm fluc}\ll L_{\rm res}$) are effectively self-averaged, larger-scale variations ($L_{\rm fluc}\sim 1$ pc) increase statistical variance by inducing local dephasing. Nevertheless, the empirical cumulative distribution function reveals that even the lower quantiles of these distributions significantly surpass the standard adiabatic case.
This resilience stems from the ”hard-driving” nature of the Floquet resonance, where the global coherence of the axion oscillation frequency $m_{a}$ exerts a persistent periodic force that facilitates mode-mixing despite environmental noise. Even when the resonance condition is fragmented by medium-scale turbulence, the collective accumulation of these imperfect resonances ensures a statistically robust conversion probability.
Standard astrophysical emission mechanisms in blazars typically yield negligible circular polarization. However, parametric resonance breaks the degeneracy between left- and right-handed circular polarization states through the parity-violating coupling. This induces a non-zero net circular polarization, whose sign and magnitude are stochastic and depend on the phase $\varphi_{a}$ of the axion background at the resonance onset.
We apply the formalism to constrain the axion–photon coupling $g_{a\gamma}$ using optical polarimetry of 3C 279. The DM density is modeled using a fuzzy DM profile, which consists of a solitonic core and an outer NFW envelope [Navarro:1995iw, Schive:2014dra] with the total mass $M_{h}\sim 10^{13}M_{\odot}$ [Ferrarese:2002ct]. The core density profile is adopted as $\rho_{c}[1+0.091(r/r_{c})^{2}]^{-8}$ with the core radius $r_{c}\simeq 1.6\,\mathrm{kpc}\left(m/10^{-22}\,\mathrm{eV}\right)^{-1}\left(M_{\rm h}/10^{9}\,M_{\odot}\right)^{-1/3}$ and density $\rho_{c}\simeq 2.9\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-3}(m/10^{-22}\,\mathrm{eV})^{2}(M_{\rm h}/10^{9}\,M_{\odot})^{4/3}$ [Schive:2014hza]. The transition radius between the two regions is taken to be $r_{t}\simeq 3r_{c}$. The concentration parameter of the NFW halo is determined using the concentration–mass relation at redshift $z=0.5$ [Dutton:2014xda]. This configuration yields negligible spatial gradients with $\mathcal{F}\lesssim 10^{-3}$ .
Figure S2: KDE-estimated probability density of induced circular polarization for various $m_{a}$ ($g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$), obtained by uniformly sampling the initial axion phase. Shaded areas denote density profiles computed using Gaussian kernels with Scott’s rule bandwidths.
For each parameter pair $(m_{a},g_{a\gamma})$, we solved the equations of motion 1,000 times, using equally spaced values between 0 and $2\pi$ for the initial axion phase $\varphi_{a}$ and assuming a $30\%$ initial linear polarization [Liodakis:2021els], then we obtained a well-resolved ensemble of induced circular polarization. In Fig. S2, the resulting probability density of the degree of circular polarization $\Pi_{C}$ is visualized using the kernel density estimation. By comparing the variance of this predicted signal with characteristic observational uncertainties, we set conservative consistency limits on the axion–photon coupling $g_{a\gamma}$ for the mass range $10^{-23}-10^{-20}\,\mathrm{eV}$ in Fig. 2.

### v2

@@
 Abstract
-Wave propagation in a coherently oscillating background is intrinsically a periodically driven problem.
-We show that photon propagation through an axion dark matter background in the presence of a magnetic field is governed by Floquet physics, distinct from conventional static or adiabatic mixing paradigms.
-Coherent photon–axion mode conversion occurs when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$) with the axion mass $m_{a}$, even far from the standard level-crossing condition $\Delta_{\gamma}\approx\Delta_{a}$.
-Crucially, this resonance disappears entirely if the axion oscillations are averaged over, and is therefore systematically missed in conventional static or adiabatic treatments.
-This driven resonance represents a unitary Floquet mode-mixing process and is fundamentally distinct from parametric instability or stimulated axion decay, preserving the axion dark matter number density.
-We develop a general Floquet framework for photon propagation in oscillating axion backgrounds, revealing that resonant mixing generates robust polarization effects during propagation.
-As an astrophysical application, we apply this mechanism to the realistic environment of the blazar 3C 279 to derive concrete constraints on the axion–photon coupling.
-While the observational manifestation depends on environmental conditions, the underlying driven mixing mechanism is generic to coherent axion dark matter, revealing a previously overlooked regime of photon–axion conversion.
+Wave propagation in periodically time-dependent media can exhibit driven mode conversion that is absent in static or adiabatic descriptions.
+We show that photon propagation through a coherent axion dark matter background provides a natural realization of such driven dynamics.
+In the presence of a magnetic field, the oscillating axion field acts as a coherent temporal drive, inducing resonant photon–axion conversion when the mismatch between their dispersion relations is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ with $n\in\mathbb{Z}$.
+This driven resonance enables efficient mixing far from the conventional level-crossing regime and disappears entirely upon time averaging, explaining why it is missed in standard treatments.
+The process constitutes a unitary mode-conversion phenomenon that preserves the axion dark matter number density and is distinct from parametric instabilities or axion decay.
+A systematic description is naturally provided by Floquet theory.
+We develop a general framework for photon propagation in oscillating axion backgrounds and show that the resulting resonant mixing leads to characteristic polarization signatures, with potential implications for astrophysical observations such as blazar polarization.
 Introduction.– The propagation of waves in time-dependent media is a central problem across physics.
-When the properties of a medium vary periodically in time, the resulting dynamics are generically governed by Floquet physics, leading to resonant mode mixing that cannot be captured by static or adiabatic approximations.
-Such driven propagation phenomena are well known in condensed matter and quantum optics [1965PhRv..138..979S, Sambe:1973cnm, 2004PhR...390....1C, Oka:2009djg, Eckardt:2016lof], yet they have received little attention in the context of particle astrophysics.
-A prominent and well-motivated realization of a coherently oscillating medium arises if dark matter (DM) [Zwicky:1933gu, Bertone:2018krk] is composed of axions [Peccei:1977hh, Peccei:1977ur, Weinberg:1977ma, Wilczek:1977pj, Cheng:1987gp] or axion-like particles [Preskill:1982cy, Abbott:1982af, Dine:1982ah, Arvanitaki:2009fg, Graham:2015ouw, Peccei:2006as, Marsh:2015xka, DiLuzio:2020wdo, Adams:2022pbo, OHare:2024nmr].
+When the properties of a medium vary periodically in time, wave propagation can exhibit driven mode mixing and resonant conversion phenomena that are absent in static or adiabatic descriptions.
+Such effects are well known in condensed matter and quantum optics [45, 41, 14, 32, 18], yet they have received little attention in the context of particle astrophysics.
+A prominent and well-motivated realization of a coherently oscillating medium arises if dark matter (DM) [50, 8] is composed of axions [34, 33, 48, 49, 13] or axion-like particles [36, 2, 16, 6, 21, 35, 27, 15, 3, 31].
 In this case, the DM forms a classical field that oscillates at a frequency set by the axion mass over macroscopic coherence scales.
 Photons propagating through this background therefore experience a medium whose optical properties are intrinsically periodic in time.
-Nevertheless, conventional treatments of photon–axion mixing [Raffelt:1987im, Raffelt:1996wa] typically average over the axion oscillations or assume an effectively static background, thereby reducing the problem to a time-independent level-crossing or adiabatic conversion scenario.
-In astrophysical environments, the photon–axion system is typically analyzed in two regimes: photon–axion conversion (the Primakoff effect) [Primakoff:1951iae, Sikivie:1983ip, Raffelt:1987im] and rotation of the polarization plane (birefringence) [Carroll:1989vb, Carroll:1998zi, Harari:1992ea].
-Conventional analyses employ adiabatic or perturbative approximations, effectively averaging over rapid axion oscillations or treating the background as quasi-static.
-In this work, we show that these assumptions miss a qualitatively novel class of resonant propagation effects.
-We demonstrate that photon–axion mixing in a coherent axion DM background is fundamentally a periodically driven problem, governed by Floquet dynamics rather than static dispersion matching.
-Our analysis reveals that the intrinsic time periodicity of the axion background acts as a parametric driver, actively inducing coherent mixing between photon and axion modes.
-This process leads to parametric resonance when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion mass as $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$).
-Consequently, resonant conversion can occur far from the conventional level-crossing resonance condition $\Delta_{\gamma}\approx\Delta_{a}$ [Pshirkov:2007st, Hook:2018iia].
-This resonance arises purely from temporal phase matching induced by the oscillating background and does not rely on spatial inhomogeneities or fine-tuned environmental profiles.
-The driven resonance is not a higher-order correction to static mixing: it relies on the coherent axion oscillation and vanishes completely once this time dependence is averaged out.
-The driven resonance identified here is fundamentally distinct from parametric instability or stimulated axion decay ($m_{a}\simeq 2\omega$) [Tkachev:2014dpa, Arza:2019nta, Caputo:2018vmy, Caputo:2018ljp].
-The propagation dynamics remain unitary, the axion background number density is conserved, and the system exhibits coherent Floquet mode mixing rather than exponential amplification.
-We develop a general Floquet framework for photon propagation in oscillating axion backgrounds and show that this mechanism leads to robust and generic consequences for photon polarization during propagation.
-While specific observational manifestations depend on environmental conditions, the underlying driven mixing phenomenon is intrinsic to coherent axion DM and represents a previously overlooked regime of photon–axion conversion.
-Our results establish a systematic framework for analyzing propagation effects in time-dependent backgrounds, revealing previously unexplored resonant regimes.
-Driven photon–axion Mixing.– Consider photon propagation in the presence of a magnetic field and a coherently oscillating axion background (see also [Espriu:2014lma, Masaki:2019ggg, McDonald:2019wou] for a similar system),
+Nevertheless, conventional treatments of photon–axion mixing [40, 39, 25] typically average over the axion oscillations or assume an effectively static background, thereby reducing the problem to a time-independent level-crossing or adiabatic conversion scenario.
+Most astrophysical studies consequently focus on birefringence [11, 12, 22] or Primakoff conversion [37, 46, 40] within quasi-static or adiabatic approximations.
+In this work, we show that these assumptions miss a qualitatively new class of resonant propagation effects.
+Photon–axion mixing in a coherent axion DM background is intrinsically a driven process, in which the temporal periodicity of the axion field acts as a coherent driver.
+Resonant mode conversion occurs when the mismatch between the photon and axion dispersion relations is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$ ($n\in\mathbb{Z}$), allowing efficient conversion even far from the conventional level-crossing condition $\Delta_{\gamma}\approx\Delta_{a}$ [38, 24].
+This resonance arises purely from temporal phase matching induced by the oscillating background and does not rely on spatial inhomogeneities or finely tuned environmental profiles.
+The effect relies on the coherence of the axion oscillations and disappears under time averaging, explaining why it is systematically missed in standard treatments.
+A systematic description of such periodically driven mixing is naturally provided by Floquet theory, originally formulated in Refs.
+[45, 41] and reviewed in Ref.
+[18].
+The driven resonance constitutes a unitary mode-conversion process, distinct from previously studied decay or instability phenomena [47, 23, 9, 10, 7], and leads to characteristic polarization signatures during propagation.
+While specific observational manifestations depend on environmental conditions, the underlying resonance is intrinsic to coherent axion DM and represents a previously unexplored regime of photon–axion mixing.
+More broadly, our results provide a general framework for wave propagation in time-dependent backgrounds and identify new resonant regimes beyond static or adiabatic descriptions.
+Driven photon–axion Mixing.– The axion may couple to electromagnetism through the operator $\mathcal{L}\supset-g_{a\gamma}aF\tilde{F}/4$, where $g_{a\gamma}$ is the photon–axion coupling with units of $\mathrm{GeV}^{-1}$, $a$ is the axion field, and $F$ is the electromagnetic field tensor.
+We consider photon propagation in the presence of a magnetic field and a coherently oscillating axion background (see also Refs.
+[19, 28, 29] for a similar system).
+The axion field is treated as a classical, homogeneous background,
 $\bar{a}(t,z)=m_{a}^{-1}\sqrt{2\rho_{\text{DM}}}\cos(m_{a}t+\varphi_{a})$ (1)
-where $m_{a}$ is the axion mass, $\rho_{\text{DM}}$ is the DM density, and $\varphi_{a}$ is the phase of the axion background.
-In an appropriate polarization basis, the propagation equation takes the form
+where $m_{a}$ is the axion mass, $\rho_{\text{DM}}$ is the DM density, and $\varphi_{a}$ an a priori unknown phase.
+The axion-photon mixing equations in the presence of an external magnetic field can be reduced to a system of first-order differential equations via the WKB approximation, assuming the beam energy $\omega$ is much larger than the axion mass $m_{a}$.
+We consider photon propagation along the $z$-axis in the presence of a transverse magnetic field $\mathbf{B}_{\mathrm{T}}$.
+In the linear polarization basis, the photon–axion system obeys the Schrödinger-like propagation equation,
 $i\partial_{z}\Psi=\left[\mathcal{H}_{0}+\mathcal{H}_{1}\right]\Psi$ (2)
-where the state vector $\Psi$ includes the two photon polarization states and the axion field.
-The photon–axion system is governed by a Hermitian mixing Hamiltonian whose coefficients contain an explicit periodic modulation at frequency $m_{a}$, induced by the axion background.
-$\mathcal{H}_{0}$ incorporates the usual photon dispersion $\Delta_{\perp}$ and $\Delta_{\parallel}$, axion mass term $\Delta_{a}$, and magnetic-field-induced mixing $\Delta_{\mathrm{B}}$, and is given by
+where the state vector $\Psi$ denotes the two photon polarization states and the axion field.
+The photon components correspond to linear polarizations oriented perpendicular and parallel to the direction of the external magnetic field.
+The dynamics is governed by a Hermitian mixing Hamiltonian with an explicit periodic time dependence induced by the oscillating axion background.
+The static part $\mathcal{H}_{0}$ describes conventional photon–axion mixing in a magnetized medium and is given by [40, 39, 25]
 $\mathcal{H}_{0}=-\begin{pmatrix}\Delta_{\perp}&0&0\\ 0&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$. (3)
-$\mathcal{H}_{1}$ encodes the axion-induced birefringence
+Throughout this work, all dispersion and mixing effects are parametrized in terms of $\Delta$-parameters following the standard photon–axion mixing notation.
+Here, $\Delta_{\perp,\parallel}$ represent the photon dispersion for the two linear polarizations, accounting for both plasma effects and vacuum birefringence.
+In the systems considered here, vacuum QED effects are included in $\Delta_{\perp,\parallel}$ but remain subdominant to the plasma contribution, $\Delta_{\perp,\parallel}\simeq-\omega_{\mathrm{pl}}^{2}/(2\omega)$.
+The effective plasma frequency $\omega_{\mathrm{pl}}=\sqrt{4\pi\alpha n_{e}/m_{e}}$ with $\alpha$ the fine-structure constant, where $n_{e}$ and $m_{e}$ denote the electron number density and mass, respectively.
+The axion mass-induced dispersion is given by $\Delta_{a}=-m_{a}^{2}/(2\omega)$, while the photon–axion mixing is governed by $\Delta_{\rm B}=g_{a\gamma}B_{\mathrm{T}}/2$.
+The axion background generates an additional, explicitly time-dependent contribution
 $\mathcal{H}_{1}(z)=-\begin{pmatrix}0&\Delta_{F}&0\\ \Delta_{F}^{*}&0&0\\ 0&0&0\end{pmatrix}$, (4)
-with
+arising from the axion-gradient-induced birefringence, analogous to but distinct from plasma Faraday rotation, with
 $\Delta_{F}=ig_{a\gamma}(\partial_{t}{\bar{a}}+\partial_{z}\bar{a})/2.$ (5)
-In the linear polarization basis, the axion-induced Faraday term is purely imaginary, reflecting its interpretation as a rotation generator.
-The Hamiltonian is manifestly Hermitian, and its only explicit time dependence arises from the oscillating axion background:
+This axion-induced Faraday term is purely imaginary, reflecting its interpretation as a rotation generator.
+While similar time-dependent terms have appeared in Refs.
+[19, 28, 29], we formalize it here as a Floquet driving parameter $\Delta_{F}$ (see Supplemental Material for the full derivation from the Lagrangian level).
+This reformulation is not merely notational; it maps the propagation problem onto a quasi-energy framework.
+Crucially, the driven contribution does not commute with the static mixing Hamiltonian, $\mathcal{H}_{0},\mathcal{H}_{1}(z)]\neq 0$.
+As a result, replacing the full time-ordered evolution operator by one generated by the averaged Hamiltonian is not a controlled approximation $\langle\mathcal{T}e^{-i\int\mathcal{H}(z)dz}\rangle\neq e^{-i\int\langle\mathcal{H}(z)\rangle dz}$.
+In particular, time averaging projects out the nonzero Floquet sectors responsible for the driven mode mixing, thereby eliminating the effect at the level of the Hamiltonian rather than suppressing it perturbatively.
+The Hamiltonian is manifestly Hermitian, and its only explicit time dependence originates from the oscillating axion background through the birefringence term,
 $\Delta_{F}=-i\Delta_{\text{DM}}\sin(m_{a}t+\varphi_{a}),$ (6)
-where the amplitude is $\Delta_{\text{DM}}\equiv g_{a\gamma}\sqrt{\rho_{\text{DM}}/2}$. Along the photon trajectory we take the eikonal propagation approximation $t\simeq z$ (with $c=1$).
-The propagation equation represents a three-level driven system with period $T=2\pi/m_{a}$.
-It is convenient to remove the explicit time dependence induced by the axion background by transforming to a co-rotating frame that follows the axion-induced phase modulation.
-This is achieved by a time-dependent unitary transformation that eliminates the oscillatory off-diagonal term at the expense of introducing harmonic couplings between propagation eigenstates.
-We perform a time-dependent unitary transformation $\Psi(z)=U(z)\Phi(z)$ to reallocate the rapid oscillations. Applying the rotation operator $U=u\oplus 1$, where $u$ acts on the photon subspace as
+where $\Delta_{\text{DM}}\equiv g_{a\gamma}\sqrt{\rho_{\text{DM}}/2}$ sets the amplitude of the axion-induced drive.
+This term oscillates at frequency $m_{a}$ and acts as a coherent temporal drive of the photon–axion system.
+Along the photon trajectory we adopt the eikonal approximation $t\simeq z$ (with $c=1$), appropriate for forward-propagating waves with a well-defined carrier frequency.
+In this limit, the temporal oscillation of the axion background translates into a periodic modulation along the propagation direction with period $T=2\pi/m_{a}$.
+The resulting propagation equation describes a three-level system subject to coherent periodic driving.
+Rather than eliminating this time dependence, it is convenient to reallocate it by transforming to a co-rotating frame that follows the axion-induced phase modulation.
+This allows the driven nature of the mixing to be made explicit while preserving the quasienergy spectrum.
+We therefore perform a time-dependent unitary transformation $\Psi(z)=U(z)\Phi(z)$ with $U=u\oplus 1$, where $u$ acts on the photon subspace as
 $u(z)=\exp\left[-i\theta(z)\sigma_{y}\right],$ (7)
-the fast oscillating sinusoidal term is absorbed into a phase-dependent rotation angle $\theta(z)=\alpha\cos\phi(z)$ with $\alpha=\Delta_{\text{DM}}/m_{a}$ and $\phi(z)=m_{a}z+\varphi_{a}$.
-This rotation is chosen such that the term generated by $U^{\dagger}i\partial_{z}U$ exactly cancels the rapidly oscillating birefringence $\Delta_{F}$.
-Assuming negligible QED birefringence with $|\Delta_{\perp}-\Delta_{\parallel}|/2\ll m_{a}$ and $\Delta_{\mathrm{DM}}$, the effective Hamiltonian becomes:
-$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}&0&\Delta_{\mathrm{B}}\sin\theta\\ 0&\Delta_{\gamma}&\Delta_{\mathrm{B}}\cos\theta\\ \Delta_{\mathrm{B}}\sin\theta&\Delta_{\mathrm{B}}\cos\theta&\Delta_{a}\end{pmatrix}$, (8)
-where $\Delta_{\gamma}=(\Delta_{\parallel}+\Delta_{\perp})/2$ is the mean photon refractive index.
-The unitary rotation is exact algebraically and does not change the quasienergy spectrum; it simply relocates the periodic dependence.
-In the regime $\Delta_{\mathrm{B}}\ll|\Delta_{\gamma}-\Delta_{a}|$, the off-resonant photon mode can be adiabatically eliminated, allowing the dynamics to be analyzed by treating each photon mode-axion pair independently within the rotating-wave approximation.
-The resulting two-level systems are governed by the couplings $G^{(1)}=\Delta_{\mathrm{B}}\sin\theta$ and $G^{(2)}=\Delta_{\mathrm{B}}\cos\theta$.
-These couplings depend on $\theta(z)$; $\sin\theta$ and $\cos\theta$ can be written as linear combinations of $e^{\pm i\theta}$.
-Using the Jacobi-Anger expansion, $e^{i\alpha\cos\phi}=\sum_{\ell}i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi}$ with the Bessel function $J_{\ell}(\alpha)$, we expand $G^{(1,2)}$ into Fourier harmonics.
-The couplings can be written as
-$G(z)=\sum_{\ell\in\mathbb{Z}}g_{\ell}e^{i\ell m_{a}z},$ (9)
-where $g_{\ell}^{(1)}$ is nonzero only for odd $\ell$, while $g_{\ell}^{(2)}$ is nonzero only for even $\ell$.
-Explicitly, the $\ell$-th harmonic coupling scales as $g_{\ell}\sim\Delta_{\mathrm{B}}J_{\ell}(\alpha)$ up to phase factors, where the Bessel function $J_{\ell}(\alpha)$ behaves roughly as an oscillatory sine or cosine that decays proportionally to $\alpha^{-1/2}$.
-In this co-rotating basis, the driven photon–axion system acquires a Floquet structure: the axion oscillation generates sidebands separated by integer multiples of $m_{a}$, and resonant mode mixing occurs when the propagation eigenvalue splitting matches one of these harmonics.
-In this basis, the axion-induced term acts as a periodic off-diagonal drive that couples propagation eigenstates separated in quasi-energy.
-Transitions between eigenstates become resonant whenever the energy mismatch $\Delta=\Delta_{\gamma}-\Delta_{a}$ is compensated by the absorption or emission of integer multiples of the driving frequency.
-We remove the free propagation phases by defining $\Phi=\operatorname{diag}\left(e^{i\Delta_{\gamma}z},e^{i\Delta_{a}z}\right)\tilde{\Phi}_{\gamma}$.
-Under this transformation, the $\ell$-th Fourier component of the coupling acquires a phase factor $e^{i(\Delta_{\gamma}-\Delta_{a}-\ell m_{a})z}$, with the corresponding detuning given by $\Delta_{\ell}=\Delta_{\gamma}-\Delta_{a}-\ell m_{a}$.
-Resonance occurs when the phase is stationary, yielding the Floquet resonance condition:
-$\Delta_{\gamma}-\Delta_{a}\approx nm_{a},\quad n\in\mathbb{Z}.$ (10)
-This condition characterizes coherent mode mixing within a time-periodic background, demonstrating that substantial conversion can occur even far from the standard level-crossing resonance $\Delta_{\gamma}=\Delta_{a}$.
-Near a given resonance $n$, keeping only the corresponding Fourier mode amounts to a rotating-wave approximation.
-The two-state system undergoes Rabi oscillations with an effective frequency $\Omega_{\rm R}=\sqrt{|g_{n}|^{2}+(\Delta_{n}/2)^{2}}$.
-The conversion probability is given by
+with $\sigma_{y}$ the second Pauli matrix.
+The rotation angle $\theta(z)=\alpha\cos\phi(z)$, with $\alpha=\Delta_{\rm DM}/m_{a}$ and $\phi(z)=m_{a}z+\varphi_{a}$, is chosen such that the contribution from $U^{\dagger}i\partial_{z}U$ exactly compensates the rapidly oscillating birefringence term $\Delta_{F}$.
+This transformation is algebraically exact and merely redistributes the periodic dependence among the Hamiltonian matrix elements.
+In the following we neglect QED vacuum birefringence, assuming $|\Delta_{\perp}-\Delta_{\parallel}|/2\ll m_{a},\Delta_{\rm DM}$.
+This condition ensures that the two photon polarization states remain effectively degenerate on the scale of the axion-induced modulation, so that the driven mixing dynamics is not suppressed by polarization dephasing.
+We define $\Delta_{\gamma}\equiv(\Delta_{\perp}+\Delta_{\parallel})/2$ as the average photon dispersion.
+Then the effective Hamiltonian becomes
+$\mathcal{H}^{\prime}(z)=-\begin{pmatrix}\Delta_{\gamma}&0&\Delta_{\mathrm{B}}\sin\theta\\ 0&\Delta_{\gamma}&\Delta_{\mathrm{B}}\cos\theta\\ \Delta_{\mathrm{B}}\sin\theta&\Delta_{\mathrm{B}}\cos\theta&\Delta_{a}\end{pmatrix}$.
+(8)
+The unitary transformation leaves the physical spectrum unchanged, but renders the periodic structure of the driven photon–axion mixing explicit, providing a natural starting point for the Floquet analysis below.
+The full system is a three-level Floquet problem and can be treated as such.
+However, in the regime $\Delta_{\rm B}\ll|\Delta_{\gamma}-\Delta_{a}|$, one of the photon propagation eigenstates is far off resonance and can be adiabatically eliminated.
+The dynamics then reduces to two independent photon–axion subsystems, each of which can be analyzed within the rotating-wave approximation.
+The effective couplings are $G^{(1)}=\Delta_{\rm B}\sin\theta(z)$ and $G^{(2)}=\Delta_{\rm B}\cos\theta(z)$, acting as coherent time-dependent drives.
+The axion-induced phase modulation is periodic, so the couplings can be expanded in a discrete set of harmonics with amplitudes controlled by Bessel functions $J_{\ell}(\alpha)$.
+Explicitly, this follows from the standard Jacobi–Anger expansion,
+$e^{i\alpha\cos\phi}=\sum_{\ell\in\mathbb{Z}}i^{\ell}J_{\ell}(\alpha)e^{i\ell\phi}.$ (9)
+Writing $\sin\theta$ and $\cos\theta$ in terms of $e^{\pm i\theta}$, the couplings can be decomposed into discrete Fourier harmonics,
+$G(z)=\sum_{\ell\in\mathbb{Z}}g_{\ell}e^{i\ell m_{a}z}.$ (10)
+Here $g_{\ell}^{(1)}$ ($g_{\ell}^{(2)}$) is nonzero only for odd (even) $\ell$, reflecting the parity structure of the two photon polarizations, and the magnitude of each harmonic scales as $g_{\ell}\sim\Delta_{\rm B}J_{\ell}(\alpha)$ up to phase factors.
+In this co-rotating basis, the driven photon–axion system naturally acquires a Floquet structure [18]: the oscillating axion background generates a ladder of sidebands separated by integer multiples of $m_{a}$, and the periodic off-diagonal couplings mediate transitions between propagation eigenstates whose quasi-energies differ by $\ell m_{a}$.
+To make this explicit, we remove the free propagation phases and move to the interaction picture by defining $\Phi=\operatorname{diag}\left(e^{i\Delta_{\gamma}z},e^{i\Delta_{a}z}\right)\tilde{\Phi}$.
+The $\ell$-th Fourier component of the coupling then acquires a phase factor $e^{i\delta_{\ell}z}$, with detuning $\delta_{\ell}=\Delta_{\gamma}-\Delta_{a}-\ell m_{a}$.
+Resonant mode mixing occurs when this phase is stationary, yielding the Floquet resonance condition
+$\Delta_{\gamma}-\Delta_{a}\approx nm_{a},\quad n\in\mathbb{Z},$ (11)
+corresponding to vanishing detuning $\delta_{\ell}$ with the identification $\ell=n$.
+This condition generalizes the conventional level-crossing criterion to a periodically driven system, demonstrating that coherent photon–axion conversion can occur even far from $\Delta_{\gamma}=\Delta_{a}$.
+The axion background thus supplies discrete energy quanta that compensate the photon–axion dispersion mismatch, enabling coherent conversion in a regime inaccessible to static treatments.
+Near a given Floquet resonance labeled by $n$, retaining only the corresponding Fourier component constitutes a rotating-wave approximation, valid when the nonresonant harmonics satisfy $|\delta_{\ell}|\gg|g_{\ell}|$ for $\ell\neq n$.
+The dynamics then reduces to a driven two-level system undergoing coherent Rabi oscillations [4, 44] with frequency
+$\Omega_{\rm R}=\sqrt{|g_{n}|^{2}+(\delta_{n}/2)^{2}}.$ (12)
+The photon–axion conversion probability takes the standard form
 $P_{\gamma\to a}\approx\frac{|g_{n}|^{2}}{\Omega_{\rm R}^{2}}\sin^{2}\big(\Omega_{\rm R}\,z\big).$ (13)
-The resonance is maintained as long as the detuning satisfies $|\Delta_{n}(z)|\lesssim 2|g_{n}|$, a regime where the peak conversion efficiency remains at least 50%.
-Note that the photon state $\gamma$ here refers to the effective photon mode under rotation.
-It can be verified that in the static limit $m_{a}\to 0$, all sidebands collapse and the system reduces exactly to conventional photon–axion mixing.
-Additionally, for small modulation amplitude $\Delta_{\mathrm{DM}}\ll m_{a}$, only the lowest harmonics contribute, and the effect reduces to a single-frequency perturbation of the standard oscillation probability.
-More detailed derivations are provided in the supplementary material.
-An equivalent description in the circular polarization basis is also discussed in the supplementary material, where the axion background appears as a periodic modulation of the refractive indices.
-Related effects of axion-induced frequency modulation have been discussed previously, primarily in the context of birefringence or modifications of the dispersion relation [McDonald:2019wou].
-In those treatments, the time-dependent phase is typically treated perturbatively or averaged over, and the associated Floquet sidebands are not explicitly identified.
-As a result, the resonant mode mixing and the corresponding Rabi oscillations discussed here were not recognized.
-We emphasize that the Floquet resonance described here is fundamentally different from parametric instability or stimulated axion decay with $m_{a}\simeq 2\omega$.
-Here, the DM number density remains conserved, with the oscillating background providing the necessary energy ”kick” to bridge the momentum mismatch between the photon and axion dispersion relations.
-The Floquet coherent mixing in a periodic background is also dynamically distinct from the phenomenon of parametric instability.
-While both involve frequency matching, the former constitutes a unitary driven two-level problem characterized by stable mode-switching.
-In contrast, parametric instability requires a non-Hermitian structure for energy pumping (typically associated with sum-frequency channels  [^1] In the context of stimulated axion decay, the instability of the photon–axion system—considering both axion DM and magnetic fields but omitting the plasma background—is explored in Ref.
-[Masaki:2019ggg], where an instability condition $\sqrt{k_{\gamma}^{2}+m_{a}^{2}}+k_{\gamma}=nm_{a}$ is empirically identified.), which is kinematically inaccessible in the difference-frequency channel.
-Consequently, our resonance analysis leads to coherent oscillations that preserve the background number density, rather than the exponential growth characteristic of unstable systems.
-More discussion of this distinction is provided in the supplementary material.
-Representative Propagation.– To illustrate the physical implications of the Floquet resonance, we consider a representative photon propagation scenario in which the photon dispersion $\Delta_{\gamma}$ varies slowly along the propagation direction, while the axion background remains coherent over the relevant length scale.
-This setup captures the essential physics of resonance crossing without requiring detailed modeling of a specific astrophysical environment.
-As the photons propagate, the slowly varying dispersion causes the mismatch $\Delta(z)=\Delta_{\gamma}(z)-\Delta_{a}$ to sweep across the Floquet resonance condition $\Delta=nm_{a}$.
-When this occurs, the driven mode coupling becomes resonant, leading to efficient photon–axion conversion.
-Importantly, this resonance depends on temporal phase matching rather than spatial level crossing, and therefore persists in the absence of sharp gradients.
-Figure 1: Evolution of the photon–axion conversion probability $P_{\gamma\to a}$ along the jet of blazar.
-The blue solid line represents the case with an axion DM background, while the orange line shows the standard adiabatic conversion without DM.
-The vertical red dashed lines indicate the locations where the resonance condition $\Delta_{\gamma}(r)-\Delta_{a}\approx nm_{a}$ is satisfied.
-Fig.
-1 shows the photon–axion conversion probability $P_{\gamma\to a}$ as a function of propagation distance for a representative choice of parameters satisfying the Floquet resonance condition.
-A pronounced enhancement occurs when $\Delta(z)$ crosses $\Delta=m_{a}$, demonstrating coherent mode mixing induced by the axion-driven modulation.
-Away from resonance, the conversion probability remains small and oscillatory, consistent with conventional photon–axion mixing.
-The resonant region acts as a phase-matched mixing layer, in which the axion oscillation continuously supplies the energy mismatch required for conversion.
-The resulting behavior is analogous to Rabi oscillations in driven two-level systems, with the conversion probability bounded by unitarity.
-We have verified that the resonance persists under moderate variations of the dispersion profile and in the presence of stochastic fluctuations, provided that the axion field remains coherent over the resonance region.
-Thus, the effect is not a fine-tuned artifact of a specific background profile.
-Detailed numerical analysis is presented in the Supplementary Material.
-A distinctive consequence of this driven mixing is the generation of polarization asymmetries between propagation modes.
-In particular, the stochastic crossing of Floquet resonances induces circular polarization with a characteristic variance set by the resonance strength.
-Since the axion-induced modulation affects the two circular polarization states with opposite signs, Floquet resonances generically break the symmetry between them.
-As a result, photon propagation through an axion DM background becomes intrinsically polarization-sensitive, even for initially unpolarized states.
-In particular, repeated or stochastic crossings of Floquet resonances lead to a net circular polarization with vanishing mean but nonzero variance.
-This stochastic circular polarization arises from coherent, unitary mode mixing and is therefore qualitatively distinct from polarization generated by absorption or scattering processes.
-The magnitude of the variance is directly controlled by the resonance strength, thereby providing a direct probe of the driven mixing mechanism.
-Importantly, this polarization signature does not rely on fine-tuned source properties or detailed environmental modeling.
-It follows directly from the existence of Floquet resonances in photon propagation through a coherent axion background.
-Constraints from the balazar 3C 279.– We validate this theoretical framework with full numerical solutions and apply it to the optical polarimetry of the blazar 3C 279 as an example.
-The parametric resonance breaks the symmetry between circular polarization states, inducing a stochastic signal in the blazar jet.
-Detailed configurations are provided in the supplementary material.
-To derive constraints on the axion–photon coupling $g_{a\gamma}$, we adopt a conservative statistical criterion: the standard deviation of the axion-induced circular polarization $\sigma_{\mathrm{std}}$ must not exceed the characteristic observational uncertainty of the blazar 3C 279 ($\sigma_{\Pi_{C}}\approx 0.30\%$ [Liodakis:2021els]).
-This ensures consistency between the predicted stochastic signal and the measured noise level.
-For a given astrophysical source, the axion phase $\varphi_{a}$ is effectively constant over observational timescales.
-However, since its value is a priori unknown, the axion-induced circular polarization behaves observationally as a stochastic contribution with zero mean.
-Even when the axion-induced circular polarization exhibits a bimodal distribution for certain masses, its sign and magnitude remain unpredictable due to the unknown axion phase, and the variance provides a conservative measure of its observational impact.
-Requiring its variance not to exceed the instrumental sensitivity therefore provides a conservative consistency bound.
-Fig.
-2 shows the illustrative constraint on parameter space, with CAST bounds [CAST:2017uph] included for comparison.
-For $m_{a}\in[10^{-23},10^{-20}]\,\mathrm{eV}$, we obtain conservative limits as $g_{a\gamma}\lesssim 2\times 10^{-11}\,\mathrm{GeV}^{-1}$.
-The distinctive band-like structure of the exclusion contour is a signature of the Floquet parametric resonance mechanism, reflecting the discrete nature of the resonance condition.
-Two features emerge: (1) The conversion efficiency, affected by the Bessel functions $J_{n}(\Delta_{\text{DM}}/m_{a})$, is maximized when the effective momentum mismatch is bridged by integer harmonics of $m_{a}$.
-The sharp dips in the exclusion limit correspond to ”sweet spots” where this resonance condition is optimally satisfied within the jet profile.
-(2) As $m_{a}$ varies, the dominant contribution shifts between different resonance orders $n$.
-Weaker constraints between bands reflect off-resonance regimes, where the system approaches the adiabatic limit, resulting in reduced circular polarization production.
-The constraints in Fig.
-2 show that axion-induced parametric resonance leads to verifiable physical consequences rather than negligible perturbative effects.
-Although the current limit region is set by the specific noise floor of 3C 279, the concrete limit validates the physical relevance of our framework.
-Figure 2: Exclusion contours on the axion–photon coupling $g_{a\gamma}$ derived from 3C 279 optical polarimetry.
+where $\gamma$ denotes the effective photon propagation eigenstate in the co-rotating basis.
+Resonant conversion is sustained as long as the detuning remains within $|\delta_{n}(z)|\lesssim 2|g_{n}|$, for which the peak conversion probability exceeds $50\%$.
+This regime defines the finite bandwidth of each Floquet resonance.
+Several consistency limits are immediate.
+In the static limit $m_{a}\to 0$, all sidebands collapse and the system reduces exactly to conventional photon–axion mixing.
+For weak modulation, $\Delta_{\rm DM}\ll m_{a}$, only the lowest harmonics contribute, and the driven resonance reduces to a small, single-frequency perturbation of the standard oscillation probability.
+Additional derivations are provided in the Supplemental Material, where we also discuss an equivalent formulation in the circular polarization basis.
+In that representation, the axion background appears as a periodic modulation of the refractive indices, providing an alternative but equivalent perspective on the driven mixing.
+Previous studies have considered axion-induced frequency modulation or birefringence [29], but typically treat the time dependence perturbatively or average over it, thereby missing the discrete Floquet sidebands and the associated resonant mode mixing.
+Finally, we stress that the driven resonance identified here—sometimes loosely termed “parametric”—corresponds to a difference-frequency Floquet resonance and is dynamically distinct from parametric instability or stimulated axion decay at $m_{a}\simeq 2\omega$ [47, 23, 9, 10, 7].
+In the present case the Hamiltonian remains Hermitian, the axion dark matter number density is conserved, and the oscillating background supplies only discrete energy quanta required to compensate the photon–axion dispersion mismatch.
+The result is bounded, unitary mode conversion characterized by coherent Rabi oscillations.
+By contrast, parametric instability requires non-Hermitian energy pumping and access to sum-frequency channels, leading to exponential growth of field amplitudes.
+Such channels are kinematically inaccessible in the difference-frequency resonance considered here.
+The driven photon–axion mixing therefore realizes a stable Floquet resonance rather than an instability, as discussed further in the Supplemental Material.
+Representative Propagation.– To illustrate the physical consequences of Floquet resonances, we consider photon propagation in a medium with a slowly varying dispersion $\Delta_{\gamma}(z)$, in the presence of a coherent axion background over the relevant length scale.
+This setup isolates the driven resonance mechanism without reliance on detailed source modeling.
+As the photon propagates, the dispersion mismatch sweeps across the Floquet resonance condition $\Delta_{\gamma}(z)-\Delta_{a}\simeq nm_{a}$, activating resonant photon–axion mode conversion.
+Unlike conventional level-crossing scenarios, this resonance is governed by temporal phase matching and therefore does not require sharp spatial gradients or adiabaticity violation.
+Figure 1: Photon–axion conversion probability $P_{\gamma\to a}$ along a representative blazar jet.
+The blue line shows the case of driven Floquet mixing with oscillating DM background, while the orange line indicates standard adiabatic conversion with static DM background.
+Vertical red dashed lines mark locations where $\Delta_{\gamma}(r)-\Delta_{a}\approx nm_{a}$.
+Figure 1 shows the resulting photon–axion conversion probability $P_{\gamma\to a}$ as a function of propagation distance for representative parameters.
+The driven Floquet case (blue) exhibits sharply enhanced conversion when the dispersion mismatch crosses a resonance band, whereas standard static mixing (orange) remains oscillatory and strongly suppressed.
+The vertical dashed lines indicate the locations where $\Delta_{\gamma}(z)-\Delta_{a}\simeq nm_{a}$.
+The resonant region acts as a phase-matched mixing layer in which the axion oscillation continuously compensates the photon–axion dispersion mismatch.
+The resulting dynamics is directly analogous to Rabi oscillations in a coherently driven two-level system, with conversion probabilities bounded by unitarity.
+We have verified that the resonance persists under moderate variations of $\Delta_{\gamma}(z)$ and in the presence of stochastic fluctuations, provided that the axion field remains coherent across the resonance region.
+The effect therefore does not rely on finely tuned environmental conditions.
+Further numerical details are presented in the Supplemental Material.
+A direct phenomenological consequence of driven resonance crossings is the generation of polarization asymmetries between propagation modes.
+Repeated or stochastic crossings induce circular polarization with vanishing mean but nonzero variance.
+Because the axion-induced modulation enters with opposite signs for the two circular polarization states, the driven mixing mechanism is intrinsically polarization sensitive, even for initially unpolarized radiation.
+The variance of the circular polarization thus provides a direct observational signature of Floquet-driven photon–axion mixing.
+Blazar 3C 279 as an illustration.– We validate this theoretical framework with full numerical solutions and demonstrate that the driven mixing mechanism can imprint characteristic signatures on photon polarization during propagation.
+As a concrete realization, we apply the formalism to optical polarimetry of the blazar 3C 279, with source-specific modeling details provided in the supplementary material.
+To illustrate the potential observational impact of the mechanism, we adopt a statistical consistency criterion based on circular polarization fluctuations.
+Specifically, we require that the standard deviation of the axion-induced circular polarization, $\sigma_{\mathrm{std}}$, does not exceed the characteristic observational uncertainty of a given source.
+For 3C 279, we take $\sigma_{\Pi_{C}}\approx 0.30\%$ [26].
+This choice reflects the fact that, while the axion field remains phase-coherent over observational timescales, its phase $\varphi_{a}$ is a priori unknown.
+As a result, the induced circular polarization does not correspond to a deterministic signal but instead manifests as a stochastic contribution with vanishing mean.
+In this situation, the variance provides a physically well-defined and observationally relevant measure of the effect.
+Applying this criterion yields an illustrative bound on the axion–photon coupling, shown in Fig.
+2 together with CAST constraints [5] for reference.
+For axion masses in the range $m_{a}\in[10^{-23},10^{-20}]\,\mathrm{eV}$, the resulting sensitivity reaches $g_{a\gamma}\lesssim 2\times 10^{-11}\,\mathrm{GeV}^{-1}$ for the adopted source parameters.
+While the numerical level of the bound depends on the specific noise properties of 3C 279, its qualitative structure is set by the underlying driven mixing dynamics.
+In particular, the banded pattern in parameter space directly reflects the discrete Floquet resonance condition associated with the periodic axion background.
+As $m_{a}$ varies, the dominant contribution shifts between different resonance orders $n$, leading to alternating regions of enhanced and suppressed conversion.
+Near resonance, the driven mixing efficiently generates circular polarization, whereas between resonances the system approaches the adiabatic regime and the effect is correspondingly reduced.
+This characteristic structure is a direct consequence of the time-dependent Floquet driving and does not rely on fine-tuned phase alignment.
+We emphasize that the bound presented here is an order-of-magnitude illustration, rather than a definitive competitive constraint.
+It demonstrates that the driven mixing mechanism can lead to observable, mechanism-specific polarization signatures at realistic sensitivity levels.
+More generally, the emergence of discrete resonance bands and stochastic circular polarization constitutes a robust phenomenological consequence of photon propagation in a coherently oscillating axion background.
+Figure 2: Illustrative bound on the axion–photon coupling $g_{a\gamma}$ derived from 3C 279 optical polarimetry.
 The shaded region indicates parameters excluded by the non-observation of excess circular polarization.
-The resulting exclusion exhibits a distinctive band structure that directly reflects the discrete Floquet resonance condition $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, providing a characteristic signature of the driven mixing mechanism.
-Conclusions.– Wave propagation in a coherently oscillating background is intrinsically a periodically driven problem.
-We have shown that photon propagation in axion DM therefore exhibits Floquet dynamics, leading to resonant photon–axion mode conversion beyond the conventional static or adiabatic mixing paradigm.
-This driven resonance arises when the dispersion mismatch between photons and axions is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, enabling coherent conversion even far from the standard level-crossing condition.
-The resulting Floquet resonance represents a distinct class of propagation phenomena.
-Unlike traditional resonant conversion, which is localized to specific spatial regions where plasma and axion masses coincide, the driven resonance identified here originates from temporal phase matching induced by the oscillating DM background.
-As a consequence, resonant mixing can persist over extended propagation distances, allowing small conversion amplitudes to accumulate coherently and imprint robust signatures on photon polarization.
-The significance of this mechanism depends on the stability of the driven system against environmental fluctuations.
-In rapidly varying media, strong gradients can disrupt the phase coherence required for Floquet mixing, suppressing the resonance.
-In sufficiently smooth environments, however, the driven propagation effect can dominate photon–axion conversion, opening a coherence-based window for probing axion dark matter that is complementary to localized, high-density resonance searches.
-More broadly, the Floquet framework developed here is not restricted to a single axion species.
-Because the resonance condition relies on the periodic structure of the background rather than its microscopic identity, analogous driven conversion phenomena may arise in multi-axion systems or more complex dark-sector environments.
-Our results therefore establish a general foundation for studying resonant wave propagation in coherently oscillating media, revealing a previously unexplored regime of dark-matter-induced photon conversion.
+The driven mixing mechanism gives rise to a distinctive band structure in parameter space, directly reflecting the discrete Floquet resonance condition $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$.
+Conclusions.– Wave propagation in a coherently oscillating background is intrinsically a periodically driven process.
+Photon propagation through axion dark matter therefore exhibits a driven resonant mode-conversion phenomenon that lies beyond the conventional static or adiabatic mixing paradigm.
+Resonant conversion occurs when the mismatch between photon and axion dispersion relations satisfies $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, allowing coherent conversion even far from the standard level-crossing condition.
+This driven resonance represents a distinct class of propagation effects.
+Unlike conventional resonant conversion, which is localized to regions where plasma and axion masses coincide, it arises from temporal phase matching induced by the oscillating background and can persist over extended propagation distances.
+Small conversion amplitudes then accumulate coherently, leading to characteristic polarization signatures.
+The effect depends on phase coherence along the propagation path.
+Rapidly varying environments suppress the resonance, while sufficiently smooth backgrounds allow driven conversion to operate effectively and, in some regimes, to dominate over static mixing contributions.
+More generally, the framework developed here applies to wave propagation in coherently oscillating backgrounds beyond the specific axion model considered.
+Because the resonance relies on temporal periodicity rather than on the microscopic origin of the oscillations, analogous driven mode conversion may arise in multi-axion or other dark-sector scenarios.
+Taken together, these results establish a systematic description of resonant wave propagation in time-dependent media and identify a previously overlooked driven regime of photon–axion mixing induced by coherent dark matter oscillations.
 Acknowledgements.
 Acknowledgments.– The authors would like to thank the computing cluster of School of Fundamental Physics and Mathematical Sciences at Hangzhou Institute for Advanced Study, UCAS for providing the computational resources used in this work.
-This work is supported by the National Key Research and Development Program of China Grant No.
-2020YFC2201502, and the National Natural Science Foundation of China under Grants No.
-12475065, 12250010, 12447131, 12447105, and 12575113.
+This work is supported by the National Natural Science Foundation of China under Grants No.
+12547110,12475065, 12250010, 12447131, 12447105, and 12575113.
 Supplementary Material for Resonant photon–axion Mixing
@@
 This Supplementary Material (SM) contains additional calculation and derivation in support of the results presented in this work.
-First, we give a full derivation of the mixing matrix $\mathcal{M}$.
+First, we give a full derivation of the Hamiltonian presented in the main text.
 Then, we present additional details regarding the derivation of the two-level systems.
@@
 I photon–axion Conversion in an Axion Background
+In this section, we provide the step-by-step derivation of the Hamiltonian presented in the main text.
 The Lagrangian for the photon field $A_{\mu}$ and axion field $a$ can be written as
@@
 For a weakly magnetized cold plasma, the dielectric tensor is isotropic, $\bm{\epsilon}=(1-\frac{\omega_{\mathrm{pl}}^{2}}{\omega^{2}})\mathbb{I}$, where $\omega_{\mathrm{pl}}=\sqrt{4\pi\alpha n_{e}/m_{e}}$ is the plasma frequency and $\omega$ is the energy of the photon–axion system.
+Here, $\alpha$ is the fine-structure constant, $n_{e}$ and $m_{e}$ denote the electron number density and mass, respectively.
 In the radiation gauge, longitudinal plasma oscillations decouple from the propagating electromagnetic modes.
@@
 […]
-$\mathcal{M}(z)=\begin{pmatrix}\Delta_{\perp}&\Delta_{F}&0\\ \Delta_{F}^{*}&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$, (S13)
-where $\Delta_{\perp}=\Delta_{\parallel}\equiv-\omega_{\mathrm{pl}}^{2}/(2\omega)$, $\Delta_{a}\equiv-m_{a}^{2}/(2\omega)$, and $\Delta_{\mathrm{B}}\equiv g_{a\gamma}B/2$.
+$\mathcal{M}(z)=\begin{pmatrix}\Delta_{\perp}&\Delta_{F}&0\\ \Delta_{F}^{*}&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$. (S13)
+Here, $\Delta_{\perp,\parallel}$ represent the photon dispersion for the two linear polarizations, accounting for both plasma effects and vacuum birefringence.
+In the systems considered here, vacuum QED effects are included in $\Delta_{\perp,\parallel}$ but remain subdominant to the plasma contribution, $\Delta_{\perp,\parallel}\simeq-\omega_{\mathrm{pl}}^{2}/(2\omega)$.
+The axion mass-induced dispersion is given by $\Delta_{a}=-m_{a}^{2}/(2\omega)$, while the photon–axion mixing is governed by $\Delta_{\rm B}=g_{a\gamma}B_{\mathrm{T}}/2$ with the transverse magnetic field $B_{\mathrm{T}}$.
 The $\Delta_{F}$ term originates from the axion-gradient-induced birefringence and is formally analogous to the Faraday rotation term.
@@
 $(i\partial_{z}+\omega+\mathcal{M})\psi=0.$ (S16)
-In the mixing matrix, we have neglected the absorption effects of photons, the contribution of the cosmic microwave background’s energy density to the dispersion relation, and the QED vacuum birefringence effect.
+In the derivation above, we have neglected the absorption effects of photons, the contribution of the cosmic microwave background’s energy density to the dispersion relation, and the QED vacuum birefringence effect.
 Assuming a small DM virial velocity $v_{\mathrm{DM}}\ll 1$, we model the background field as
@@
 $\mathcal{H}_{C}^{\prime}=-\begin{pmatrix}\Delta_{+}&\delta\Delta_{\gamma}&\frac{\Delta_{\mathrm{B}}}{\sqrt{2}}\\ \delta\Delta_{\gamma}&\Delta_{-}&\frac{i\Delta_{\mathrm{B}}}{\sqrt{2}}\\ \frac{\Delta_{\mathrm{B}}}{\sqrt{2}}&-\frac{i\Delta_{\mathrm{B}}}{\sqrt{2}}&\Delta_{a}\end{pmatrix}$ (S41)
-where $\Delta_{\pm}=\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi$. Neglecting the QED birefringence effect, the evolution for the photon state can be formally written as
+where $\Delta_{\pm}=\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi$ represent the photon dispersion for the two circular polarizations.
+Neglecting the QED birefringence effect, the evolution for the photon state can be formally written as
 $\Phi_{\pm}$ $\simeq\exp\left[-i\int^{z}dz^{\prime}(\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi(z^{\prime}))\right]\tilde{\Phi}_{\pm}$ (S42)
@@
 Repeating the Fourier harmonic expansion as above, one would also obtain the same resonance condition and Rabi oscillations.
-Although axion-induced frequency modulation has been noted in earlier studies [McDonald:2019wou], it is often incorporated as a slowly varying phase or an effective modification of the refractive index.
+It has been noted that coherent axion background induces a periodic frequency modulation of photon propagation in Ref.
+[29].
+However, in that work the modulation is treated as a phase modulation of a fixed propagation eigenmode, and physical observables are obtained after averaging over the axion oscillation.
+As a result, the associated sidebands are not promoted to independent dynamical degrees of freedom.
+This procedure effectively restricts the dynamics to a single Floquet sector and is equivalent to replacing the time-ordered evolution operator by one generated by an averaged Hamiltonian.
 Such treatments effectively average over the harmonic structure of the modulation and therefore do not capture the Floquet sidebands responsible for resonant mode mixing.
@@
 When $m_{a}\simeq 2\omega$, the system enters an instability band where the field amplitude grows exponentially.
+[^1] The instability of the photon–axion system—considering both axion DM and magnetic fields but omitting the plasma background—is explored in Ref.
+[28], where an instability condition $\sqrt{k_{\gamma}^{2}+m_{a}^{2}}+k_{\gamma}=nm_{a}$ is identified.
 This corresponds to the stimulated decay of the axion field into photon pairs, a genuine parametric instability where the background acts as an energy pump.
@@
 We model the relativistic jet of blazar 3C 279, focusing on the unpolarized optical R-band photons ($E_{\gamma}\approx 1.9\,\mathrm{eV}$) with a representative parameter set: an axion mass $m_{a}=10^{-22}\,\mathrm{eV}$ and a coupling constant $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$.
-The radiation zone, located $r_{0}=0.5\,\mathrm{pc}$ from the central engine, is characterized by a magnetic field $B_{0}=0.5\,\mathrm{G}$ and electron density $n_{e,0}=5\times 10^{3}\,\mathrm{cm}^{-3}$ [A:2025wjt].
-The jet is modeled as a conical outflow with a Doppler factor $\delta\simeq 13$ [A:2025wjt], assuming scaling profiles $B(r)\propto r^{-1}$ and $n_{e}(r)\propto r^{-2}$ over a total length $L_{\rm jet}=1\,\mathrm{kpc}$.
-To capture the fine-grained evolution of the mixing process, the calculation is performed over a domain size ranging from $5\times 10^{-7}\,\mathrm{pc}$ to $10^{-3}\,\mathrm{pc}$.
-To isolate and illustrate the resonance structures dictated by the plasma profile, we adopt a uniform DM density $\rho_{\rm DM}=1\,\mathrm{GeV\cdot cm^{-3}}$ ($2.63\times 10^{7}\,M_{\odot}\mathrm{\cdot kpc^{-3}}$).
-The numerical results of the photon–axion conversion probability in Fig.
-1 demonstrate resonant conversion exactly at the locations predicted by Eq.
+The radiation zone, located $r_{0}=0.5\,\mathrm{pc}$ from the central engine, is characterized by a magnetic field $B_{0}=0.5\,\mathrm{G}$ and electron density $n_{e,0}=5\times 10^{3}\,\mathrm{cm}^{-3}$ [1].
+The jet is modeled as a conical outflow with a Doppler factor $\delta\simeq 13$ [1], assuming scaling profiles $B(r)\propto r^{-1}$ and $n_{e}(r)\propto r^{-2}$ over a total length $L_{\rm jet}=1\,\mathrm{kpc}$.
+To capture the fine-grained evolution of the mixing process, the calculation is performed over domain sizes ranging from $5\times 10^{-7}\,\mathrm{pc}$ to $10^{-3}\,\mathrm{pc}$.
+To isolate and illustrate the resonance structures dictated by the plasma profile, we adopt a uniform DM density $\rho_{\rm DM}=1\,\mathrm{GeV\cdot cm^{-3}}$ ($2.63\times 10^{7}\,M_{\odot}\mathrm{\cdot kpc^{-3}}$), and obtain the photon–axion conversion probability shown in Fig.
+1.
+Fig.
+1 demonstrates resonant conversion exactly at the locations predicted by Eq.
 (11), while the behavior reduces to the adiabatic limit away from resonance.
@@
 Thus, while a realistic non-uniform DM halo would modulate the amplitude of each ”step” according to local $\rho_{\rm DM}$ and $B$ values, the qualitative structure of multiple resonances remains unchanged.
-The spatial width of the parametric resonance $L_{\text{res}}$ can be estimated as $\left|4g_{n}/(\nabla_{r}\Delta_{n})\right|_{r_{\text{res}}}$.
+The spatial width of the Floquet resonance $L_{\text{res}}$ can be estimated as $\left|4g_{n}/(\nabla_{r}\delta_{n})\right|_{r_{\text{res}}}$.
 For the jet model considered here, the resonance width is approximately $2(nm_{a})^{-1}g_{a\gamma}B_{0}J_{n}(\alpha)r_{0}$.
@@
 Given that $L_{\text{res}}$ is two orders of magnitude smaller than the total jet length ($\sim\mathrm{kpc}$), plasma density variations within each resonance region are negligible.
-This separation of scales ensures that the parametric resonance remains robust even in highly inhomogeneous jet environments, where the global density gradient merely serves to sequentially satisfy discrete resonance conditions.
-The parametric resonance relies on the coherent oscillation of the axion field.
+This separation of scales ensures that the resonance remains robust even in highly inhomogeneous jet environments, where the global density gradient merely serves to sequentially satisfy discrete resonance conditions.
+The Floquet resonance relies on the coherent oscillation of the axion field.
 The validity of this assumption depends on the coherence time of the DM field, $\tau_{\mathrm{coh}}\sim(m_{a}v^{2})^{-1}$, compared to the photon’s transit time through the resonance region, $t_{\mathrm{cross}}\sim L_{\mathrm{res}}$.
@@
 Moreover, the fractional frequency spread $\delta m_{a}/m_{a}\sim 10^{-6}$ induces a negligible broadening of the resonance layer compared to its physical width determined by the jet geometry and coupling strength.
-Figure S1: ECDF of the photon-to-axion conversion probability $P_{\gamma\to a}$.
-The color coding indicates various turbulence correlation lengths $L_{fluc}$ for electron density (from $10^{-3}$ to $1\text{ pc}$), while solid and dashed lines distinguish between models with and without DM background, respectively.
-Each curve is derived from a sample size of $N=1000$.
-Note that the x-axis is presented in log-scale.
-To assess the robustness of the resonance mechanism, we introduce 50% stochastic electron density fluctuations across turbulence scales ranging from $L_{\rm fluc}=10^{-3}$ to $1$ pc.
+(a) Fluctuations in electron density
+(b) Fluctuations in magnetic field
+Figure S1: Empirical cumulative distribution function (ECDF) of the photon-to-axion conversion probability $P_{\gamma\to a}$.
+The colors represent different fluctuation correlation lengths $L_{\rm fluc}$, ranging from $10^{-3}$ to 11 pc.
+Panels (a) and (b) correspond to stochastic fluctuations in the electron density and magnetic field, respectively.
+Solid and dashed lines distinguish between driven Floquet mixing with oscillating DM background and adiabatic conversion with static DM background.
+Each curve is generated from $N=1000$ random realizations.
+Note that the $x$-axis is on a logarithmic scale.
+The phenomenological impact of this mechanism depends on the degree to which phase coherence is maintained along the propagation path.
+Strong gradients or rapid environmental variations can disrupt the driven mixing and suppress the resonance, while sufficiently smooth backgrounds allow the periodically driven dynamics to operate effectively.
+In such regimes, the driven contribution to photon–axion mixing can be comparable to, or exceed, that expected from static conversion alone.
+To assess the robustness of the resonance mechanism, we investigate the effects of stochastic fluctuations in the electron density and the magnetic field separately.
+The propagation path is partitioned into discrete domains of length $L_{\rm fluc}$.
+For each scenario, we introduce a random bias by scaling the baseline model value by a coefficient $\xi$, uniformly sampled from $[0.5,1.5]$, thereby representing $50\%$ amplitude fluctuations.
 As shown in Fig.
-S1, the mean conversion probability in the DM background remains consistently 1-2 orders of magnitude higher than the standard baseline without DM, regardless of the correlation length.
+S1, even with these 50% amplitude variations, the mean conversion probability under Floquet-driven mixing remains consistently 1–2 orders of magnitude higher than in the standard static scenario across all considered correlation lengths.
 While small-scale fluctuations ($L_{\rm fluc}\ll L_{\rm res}$) are effectively self-averaged, larger-scale variations ($L_{\rm fluc}\sim 1$ pc) increase statistical variance by inducing local dephasing.
-Nevertheless, the empirical cumulative distribution function reveals that even the lower quantiles of these distributions significantly surpass the standard adiabatic case.
+Nevertheless, the empirical cumulative distribution function (ECDF) reveals that even the lower quantiles of these distributions significantly surpass the standard adiabatic case.
 This resilience stems from the ”hard-driving” nature of the Floquet resonance, where the global coherence of the axion oscillation frequency $m_{a}$ exerts a persistent periodic force that facilitates mode-mixing despite environmental noise.
@@
 Standard astrophysical emission mechanisms in blazars typically yield negligible circular polarization.
-However, parametric resonance breaks the degeneracy between left- and right-handed circular polarization states through the parity-violating coupling.
+However, the Floquet resonance breaks the degeneracy between left- and right-handed circular polarization states through the parity-violating coupling.
 This induces a non-zero net circular polarization, whose sign and magnitude are stochastic and depend on the phase $\varphi_{a}$ of the axion background at the resonance onset.
+This stochastic circular polarization arises from coherent, unitary mode mixing and is therefore qualitatively distinct from polarization generated by absorption or scattering processes.
+The magnitude of the variance is directly controlled by the resonance strength, thereby providing a direct probe of the driven mixing mechanism.
+Importantly, this polarization signature does not rely on fine-tuned source properties or detailed environmental modeling.
+It follows directly from the existence of Floquet resonances in photon propagation through a coherent axion background.
 We apply the formalism to constrain the axion–photon coupling $g_{a\gamma}$ using optical polarimetry of 3C 279.
-The DM density is modeled using a fuzzy DM profile, which consists of a solitonic core and an outer NFW envelope [Navarro:1995iw, Schive:2014dra] with the total mass $M_{h}\sim 10^{13}M_{\odot}$ [Ferrarese:2002ct].
-The core density profile is adopted as $\rho_{c}[1+0.091(r/r_{c})^{2}]^{-8}$ with the core radius $r_{c}\simeq 1.6\,\mathrm{kpc}\left(m/10^{-22}\,\mathrm{eV}\right)^{-1}\left(M_{\rm h}/10^{9}\,M_{\odot}\right)^{-1/3}$ and density $\rho_{c}\simeq 2.9\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-3}(m/10^{-22}\,\mathrm{eV})^{2}(M_{\rm h}/10^{9}\,M_{\odot})^{4/3}$ [Schive:2014hza].
+The DM density is modeled using a fuzzy DM profile, which consists of a solitonic core and an outer NFW envelope [30, 42] with the total mass $M_{h}\sim 10^{13}M_{\odot}$ [20].
+The core density profile is adopted as $\rho_{c}[1+0.091(r/r_{c})^{2}]^{-8}$ with the core radius $r_{c}\simeq 1.6\,\mathrm{kpc}\left(m/10^{-22}\,\mathrm{eV}\right)^{-1}\left(M_{\rm h}/10^{9}\,M_{\odot}\right)^{-1/3}$ and density $\rho_{c}\simeq 2.9\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-3}(m/10^{-22}\,\mathrm{eV})^{2}(M_{\rm h}/10^{9}\,M_{\odot})^{4/3}$ [43].
 The transition radius between the two regions is taken to be $r_{t}\simeq 3r_{c}$.
-The concentration parameter of the NFW halo is determined using the concentration–mass relation at redshift $z=0.5$ [Dutton:2014xda].
+The concentration parameter of the NFW halo is determined using the concentration–mass relation at redshift $z=0.5$ [17].
 This configuration yields negligible spatial gradients with $\mathcal{F}\lesssim 10^{-3}$ .
-Figure S2: KDE-estimated probability density of induced circular polarization for various $m_{a}$ ($g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$), obtained by uniformly sampling the initial axion phase.
-Shaded areas denote density profiles computed using Gaussian kernels with Scott’s rule bandwidths.
-For each parameter pair $(m_{a},g_{a\gamma})$, we solved the equations of motion 1,000 times, using equally spaced values between 0 and $2\pi$ for the initial axion phase $\varphi_{a}$ and assuming a $30\%$ initial linear polarization [Liodakis:2021els], then we obtained a well-resolved ensemble of induced circular polarization.
+Figure S2: Probability density of induced circular polarization for various axion masses $m_{a}$ (assuming $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$), estimated via Kernel Density Estimation (KDE) by uniformly sampling the axion phase $\varphi_{a}\in[0,2\pi]$.
+The shaded profiles are generated using Gaussian kernels with Scott’s rule bandwidths.
+The transition from single-peaked to double-peaked distributions illustrates the varying sensitivity of the conversion process to the axion phase across different mass regimes.
+For each parameter pair $(m_{a},g_{a\gamma})$, we solved the equations of motion 1,000 times, using equally spaced values between 0 and $2\pi$ for the axion phase $\varphi_{a}$ and assuming a $30\%$ initial linear polarization [26], then we obtained a well-resolved ensemble of induced circular polarization.
 In Fig.
 S2, the resulting probability density of the degree of circular polarization $\Pi_{C}$ is visualized using the kernel density estimation.
-By comparing the variance of this predicted signal with characteristic observational uncertainties, we set conservative consistency limits on the axion–photon coupling $g_{a\gamma}$ for the mass range $10^{-23}-10^{-20}\,\mathrm{eV}$ in Fig.
-2.
+To derive constraints on the axion–photon coupling $g_{a\gamma}$, we adopt a source-agnostic statistical criterion: the standard deviation of the axion-induced circular polarization $\sigma_{\mathrm{std}}$ must not exceed the characteristic observational uncertainty of a given source.
+For the case of 3C 279, we take $\sigma_{\Pi_{C}}\approx 0.30\%$ [26].
+This criterion ensures consistency between the predicted stochastic signal and the measured noise level without assuming a deterministic phase alignment.
+For any astrophysical source, the axion phase $\varphi_{a}$ is effectively constant over observational timescales.
+However, since its value is a priori unknown, the axion-induced circular polarization manifests as a stochastic contribution with vanishing mean.
+Even when the axion-induced circular polarization exhibits a bimodal distribution for certain masses, its sign and magnitude remain observationally unpredictable due to the unknown axion phase, making the variance a physically meaningful measure of its impact.
+Requiring its variance not to exceed the instrumental sensitivity therefore provides a consistency bound that is insensitive to detailed phase realizations.
+Figure 2 shows the resulting illustrative bound in parameter space, with CAST limits [5] included for scale comparison.
+For $m_{a}\in[10^{-23},10^{-20}]\,\mathrm{eV}$, this procedure yields limits at the level of $g_{a\gamma}\lesssim 2\times 10^{-11}\,\mathrm{GeV}^{-1}$.
+The distinctive band-like structure of the resulting contour is a characteristic fingerprint of the Floquet-driven mode mixing mechanism, requiring driven, time-dependent mixing and the associated discrete resonance structure.
+Two generic features emerge from this mechanism: (1) The conversion efficiency, affected by the Bessel functions $J_{n}(\Delta_{\text{DM}}/m_{a})$, is maximized when the effective momentum mismatch is bridged by integer harmonics of $m_{a}$, a direct consequence of the periodic Floquet driving.
+The sharp dips in the consistency bound correspond to resonance “sweet spots” where this condition is optimally satisfied, largely independent of the detailed jet profile.
+(2) As $m_{a}$ varies, the dominant contribution shifts discretely between different resonance orders $n$, producing the characteristic banded pattern.
+Weaker constraints between bands reflect off-resonance regimes, where the system approaches the adiabatic limit and the driven mixing becomes inefficient, resulting in suppressed circular polarization production.
+Although the current limit region is normalized to the specific noise floor of 3C 279, the qualitative structure of the constraint reflects a phenomenological consequence of the underlying mechanism, thereby illustrating the physical relevance of the framework.